### v7

# Entropy Collapse in Mobile Sensors: The Hidden Risks of Sensor-Based Security

Abstract
Mobile sensor data has been proposed for security-critical applications such as device pairing, proximity detection, and continuous authentication. However, the foundational premise that these signals provide sufficient entropy remains under-explored. In this work, we systematically analyse the entropy of mobile sensor data across four diverse datasets spanning multiple contexts. Our findings reveal pervasive biases, with single-sensor mean min-entropy values ranging from 3.408–4.483 bits ($\sigma$=1.018–1.574) despite Shannon entropy being several multiples higher, showing a significant collapse between average- to worst-case settings. We further demonstrate that correlations between sensor modalities reduce the worst-case entropy of using multiple sensors by up to $\approx$75compared to average-case Shannon entropy. This brings joint min-entropy well below 10 bits in many cases and, in the best case, yielding only $\approx$24 bits of min-entropy when combining 20 sensor modalities. Our results demonstrate that adversaries may feasibly predict sensor signals through an exhaustive, brute-force exploration of the entire measurement space. Our work also calls into question the widely held assumption that adding more sensors inherently yields higher security, and we strongly urge caution when relying on mobile sensor data for security applications.
keywords:
Entropy , Sensors , Mobile security
1 Introduction
Modern mobile devices are equipped with various embedded sensors—accelerometers, gyroscopes, magnetometers, and others—that capture continuous motion and environmental data at fine temporal granularity. This sensor data has enabled myriad applications from activity recognition to context-aware computing. More recently, research has proposed leveraging these signals for security-critical tasks such as cryptographic key generation, zero-interaction device pairing, and continuous authentication. A crucial yet under-explored assumption underpins such designs: sensor data provides sufficient unpredictability to thwart adversarial inference. Traditional “shake-to-pair” protocols [1] rely on motion patterns to establish secure communication between co-located devices, while other methods have incorporated ambient phenomena, such as characteristics of magnetic fields and thermal fluctuations, to mitigate relay attacks [2, 3, 4, 5, 6] and reduce user authentication prompts [7, 8, 9, 10, 11].
Despite these advancements, fundamental issues remain: multi-modal sensing is often advocated to counter sensor-specific weaknesses [2, 4, 12, 13], but the quantitative security benefits of combining multiple sensors has not been rigorously evaluated. Many existing studies rely on heuristic assessments or machine learning classifiers, e.g. [4, 13, 14, 2, 12], that do not address critical security questions. That is, firstly, how much entropy do sensors truly provide? And, secondly, to what extent do multi-modal sensor combinations provide security gains? Understanding the underlying entropy is important: even if different sensors are combined, fused or otherwise transformed, it does not fundamentally improve the quantity of entropy, or unpredictability, inherent in such signals. This paper investigates those concerns.
Our analysis reveals systemic limitations: commodity sensors exhibit significant biases of between 3.408–4.483 bits of min-entropy (5.584–9.266 bits of Shannon entropy on average). In this paper, we analyse 25 different sensors compared to a far smaller number explored in related work, i.e. [15] (1 sensor), [16] (10), [7] (2), and [17] (3). Furthermore, to the best of our knowledge, we also present the first multi-modal entropy analysis at this scale. We find that, while multi-modal sensor usage confers some benefits, non-uniform distributions and inter-sensor correlations significantly reduce the worst-case min-entropy by $\approx$40–75compared to average-case Shannon entropy. These findings challenge the notion that increasing the number of sensors reliably strengthens security, and it underscores the inadequacy of using sensor data as a dependable entropy source. Our contributions are as follows:
  • 1.
    We introduce the first systematic approach to evaluating sensor entropy across such a comprehensive range of modalities and datasets using various entropy metrics (max, Shannon, collision, and min-entropy).
  • 2.
    We empirically demonstrate how inter-sensor correlations and biases erode entropy, casting doubt on the proposition that using multiple sensors adds substantially to security.
  • 3.
    We show how the collapse in worst-case entropy opens the door to attacks that exhaustively enumerate, or brute force, the space of possible measurements. This gives rise to fundamental security risks to schemes that rely on signals from single and multiple mobile sensors.
  • 4.
    Ultimately, we do not recommend relying on commodity sensors as a source of unpredictability for security-critical applications, both on a single- and multiple-sensor basis.
The rest of this paper is organised in the following way: §2 discusses sensor-based security mechanisms and established entropy metrics. §3 explains our experiment design for entropy estimation, including the threat model and dataset selection. §4 presents empirical results of our analyses and §5 discusses the implications for system design. We conclude in §6 with recommendations for further work. Our analysis work is released publicly to foster future research. [^2] https://github.com/cgshep/entropy-collapse-mobile-sensors
2 Background
This section discusses sensor-based applications, critiques existing approaches, and formalises the entropy metrics underpinning our analysis.
2.1 Mobile Sensor-based Security Applications
One major security application of mobile sensors is proximity detection, particularly for mitigating man-in-the-middle and relay attacks on mobile devices [18, 13, 14, 4, 2]. Mehrnezhad et al.[13] introduced a technique that uses accelerometer readings to verify that an NFC payment instrument and terminal are physically tapped together. The authors posit that “physical tapping causes random but correlated vibrations at both devices, which are hard to forge (or reproduce)” (p.1, [13]). The work reports an equal error rate (EER) of 17.65% using a machine learning-based approach. Gurulian et al. [14] also explored the use of shared vibration patterns between users using unique vibration patterns generated by one device. Shrestha et al. [2] explore four environmental modalities, including temperature and humidity, for proximity detection. In this work, single sensors yields 0.733–0.881 F1-score, while combining multiple increases the overall performance to 0.913–0.957. Mobile sensors have also found utility in tackling the longstanding problem of continuous authentication [19, 20, 21, 8, 9, 22, 10, 11], where sensor data is used to authenticate users passively without explicit interaction (with up to 99% accuracy claimed in some work [20]). Sensors have also been used to underpin the security of novel device pairing schemes, whether as a primary [23] or second line of authentication [1]. These schemes generally rely on detecting similar motion patterns between two devices using joint accelerometer and gyroscope sensor measurements in order to provide evidence of co-location.
The notion of ‘hardness’ is typically inferred through model evaluation metrics. A general approach follows one whereby sensor data is collected from $N$ users from which various features are extracted in the time or frequency domain (e.g. cross-correlation, spectral energy, Hamming, Euclidean and mean-absolute distances) [4, 13, 3, 2]. Features are then classified using simple threshold-based or supervised classification models, e.g. Support Vector Machines (SVM) and Random Forests. Calculating false positive (FPR) and negative (FNR) rates [1], precision and recall [8], EERs [14, 13], Receiver-Operator Curves (ROC) [11], and accuracy [11] are used to evaluate the model with respect to distinguishing between legitimate and illegitimate samples. A system is deemed to be effective if the model can discriminate between such samples with low error. (User studies have also been employed to evaluate the effectiveness of sensor-based authentication systems [21, 1].)
Some proposals have attempted to assess the entropy of sensor signals within the context of a security mechanism, but this represents a minority of work in the literature. T2Pair by Li et al. [24], a zero-interaction pairing protocol, is found to have 32.3–38.5 bits of Shannon entropy; a refinement by Wu et al. [25] reported 51–54 bits. Even in the best cases, this is low relative to modern cryptographic standards. [^3] See AIS 20/31 [26]: FCS_RNG.1 specifies $\geq$240 and $\geq$250 bits of min- and Shannon entropy respectively for the effective internal state of a random number generator. We also point to work that has questioned the utility of mobile sensors in time-critical domains. Markantonakis et al. [4], building on Gurulian et al. [5] and Shepherd et al. [6], presented a reproducibility study of mobile sensors when deployed under a 500ms time constraint for NFC-based transactions as specified by the EMV payment protocol. Here, 0.179–0.246 EER was reported depending on the given sensor combination. Sensors were thus deemed unsuitable for proximity and relay attack detection without posing usability and security issues in practice.
2.2 Sensor Entropy Analyses
In earlier work, Voris et al. [15] investigated accelerometers as true random number generators (TRNGs) on a WISP RFID tag and Nokia N97 phone. The authors find that min-entropy—defined in §2.3—is proportional to the motion applied to the device, with stationary movement having the lowest min-entropy. Intrinsic noise from the sensor’s circuitry and seismic noise, and the sampling rate of its analog-to-digital converter (ADC), are considered significant influences on entropy generation. Min-entropy values of 3.1–11.4 bits were measured depending on the movement of the accelerometer. Lv et al. [17] analysed three mobile sensors on an undisclosed Xiaomi Redmi smartphone: a triaxial accelerometer, gyroscope, and magnetometer. Min-entropy values of 0.593–5.876 are reported, depending on the modality and the entropy estimation method. Krhovják et al. [7] examined the entropy of image and audio data collected from mobile phone cameras and microphones respectively. Using Nokia N73 and E-Ten X500 and M700 phones, Shannon entropies of 2.9 (microphone) and 2.408–5.376 (camera) are reported, with min-entropy of 0.5 (microphone) and 0.754–3.928 (camera). Hennebert et al. [16] presented an analysis of 10 sensors on two wireless sensor monitors: a TI eZ430-RF2500 and a Zolertia Z1. A single-sensor analysis is presented, yielding min-entropy values of 0–7.85; motion sensors, e.g. accelerometer and vibration sensors, produced the highest entropy.
Sensors have also been suggested as entropy sources in low-cost RNG designs for mobile devices. Suciu et al. [27] proposed using a phone’s GPS module along with its accelerometer, gyroscope and orientation sensors. Using data from an HTC Google Nexus One, the approach passes the tests established in the NIST SP 800-22 [28] suite, but no precise entropy values were presented. Wallace et al. [29] explored an RNG design using the accelerometer, gyroscope, microphone, WiFi, GPS and camera data as randomness sources. Results are presented from a non-standard entropy evaluation using 37 Android devices. Sensors within existing work serve as inherently opportunistic entropy sources, i.e. in contrast with dedicated TRNGs using ring oscillators, Johnson-Nyquist thermal noise, and quantum phenomena (e.g. see [30]). Mobile sensors depend heavily on user behaviour and environmental, which could result in biases and correlations that are absent in controlled entropy sources. Existing sensor-based mechanisms largely overlook these dynamics, relying principally on heuristic or model evaluation metrics [8, 11]. Such approaches do not account well for skewed distributions and other biases in the underlying data that affect predictability. Contrast this with typical measures used in the area of randomness testing and authentication [31, 32, 30, 33, 26]. For instance, NIST SP800-90B [33] and AIS 20/31 [26] recommend the use of min-entropy to estimate ‘worst-case’ unpredictability of a given source. Our study bridges the gap by systematically evaluating entropy across modalities and datasets.
2.3 Definitions
We use the following definitions and notation throughout this work.
Definition 2.1 (Rényi Entropy).
Let $X$ be a discrete random variable taking values in a set $\mathcal{X}$ with probability mass function $p(x)$. The Rényi entropy of order $\alpha$ ($\alpha>0$, $\alpha\neq 1$) is defined as
$H_{\alpha}(X)\,=\,\frac{1}{1-\alpha}\,\log\!\Biggl{(}\sum_{x\in\mathcal{X}}p(x)^{\alpha}\Biggr{)}.$ (1)
We draw attention to four special cases of $\alpha$ that are widely used in the literature. Firstly, the Hartley (Max) Entropy, given in Eq. 2, is the logarithm of the number of possible outcomes that have non-zero probability; it serves effectively as an upper bound.
$H_{0}(X)\equiv\lim_{\alpha\to 0}H_{\alpha}(X)\,=\,\log\Bigl{|}\bigl{\{}\,x\in\mathcal{X}:p(x)>0\bigr{\}}\Bigr{|}.$ (2)
Second is the Shannon Entropy (Eq. 3), which corresponds to the classical definition of entropy in information theory, and is the limit of $H_{\alpha}$ as $\alpha\to 1$.
$H_{1}(X)\equiv\lim_{\alpha\to 1}H_{\alpha}(X)\,=\,-\sum_{x\in\mathcal{X}}p(x)\,\log p(x).$ (3)
Another case is the Collision Entropy ($\alpha=2$), quantifying the probability of “collisions” of multiple draws from $X$. This is given in Eq. 4.
$H_{2}(X)\,=\,-\log\Bigl{(}\sum_{x\in\mathcal{X}}p(x)^{2}\Bigr{)}.$ (4)
$H_{1}$ provides an average-case measure of uncertainty. It takes into account the entire distribution of outcomes; however, an adversary may only need to guess the most likely event to gain an advantage. Min-entropy is thus used as a conservative, worse-case metric, accounting for the least favorable distribution of outcomes. This is the Min-Entropy (Eq. 5), i.e. the value in the limit $\alpha\to\infty$.
$H_{\infty}(X)\equiv\lim_{\alpha\to\infty}H_{\alpha}(X)\,=\,-\log\Bigl{(}\max_{x\in\mathcal{X}}p(x)\Bigr{)}.$ (5)
Note that $H_{\alpha}$ is a non-increasing function of $\alpha$, i.e. $H_{\infty}(X)\leq H_{2}(X)\leq H_{1}(X)\leq H_{0}(X)$. $H_{\infty}(X)$ focuses on the single most likely outcome, providing a strictly tighter (and generally minimum) bound on uncertainty. We observe that min-entropy ensures that even the most skewed probability distributions still meet the required security guarantees; indeed, it is a recommended method for assessing entropy sources in NIST SP800-90B [33] and AIS 20/31 [26]. We also rely on joint entropy for assessing the entropy of multiple random variables.
Definition 2.2 (Joint Rényi Entropy).
Let $X_{1},X_{2},\dots,X_{n}$ be discrete random variables that jointly take values in $\mathcal{X}_{1}\times\mathcal{X}_{2}\times\cdots\times\mathcal{X}_{n}$, with joint probability mass function $p(x_{1},x_{2},\dots,x_{n})$. The Rényi entropy of order $\alpha$ ($\alpha>0,\alpha\neq 1$) for these $n$ variables is defined as the following, where the sum is taken over all $(x_{1},\dots,x_{n})$ in $\mathcal{X}_{1}\times\mathcal{X}_{2}\times\cdots\times\mathcal{X}_{n}$:
$H_{\alpha}(X_{1},X_{2},\dots,X_{n})\,=\,\frac{1}{1-\alpha}\,\log\Biggl{(}\sum_{(x_{1},\dots,x_{n})}p(x_{1},\dots,x_{n})^{\alpha}\Biggr{)}$ (6)
In the limit $\alpha\to 1$, then $H_{\alpha}(X_{1},X_{2},\dots,X_{n})$ converges to the classical joint Shannon entropy of these $n$ variables. To support the later discussion on Chow-Liu trees—our approach to joint entropy estimation—we also define the Kullback-Leibler divergence.
Definition 2.3 (Kullback-Leibler (KL) Divergence).
Given a true probability distribution $P(x)$ of a random variable and an approximate or reference distribution $Q(x)$, the KL divergence is defined as follows:
$D_{KL}(P\;||\;Q)=\sum_{x\in\mathcal{X}}P(x)\log\Big{(}\frac{P(x)}{Q(x)}\Big{)}$ (7)
3 Experiment Design
3.1 Threat Model and Assumptions
We consider an adversary aiming to compromise sensor-based security schemes (e.g., key generation, proximity detection, and continuous authentication). The adversary may gather extensive statistical data about how smartphone sensors behave in everyday usage; for instance, from widely available open datasets. We assume the attacker focuses on predicting or guessing the sensor outputs by prioritising the most probable values first, exploiting any biases in the distribution of sensor measurements. As such, they may resort to exhaustive enumeration of the measurement space if the target source’s entropy is low enough. We exclude capabilities such as fault injection and other hardware attacks (see [34]). While those could further reduce the effective entropy space—say, by inducing errors in the output values of sensing hardware—we regard them as outside the scope of this work. This threat model thus represents an adversary who can capitalise on statistical biases in sensor data without directly comprising the device physically. Our goal is to evaluate whether sensor data distributions—even aggregated from diverse users—offer sufficient entropy to resist attacks that search the space of sensor measurement values informed by their statistical properties.
3.2 High-level Methodology
We aim to determine the global (i.e., population-level) entropy characteristics of various mobile sensors under ordinary usage conditions, rather than focusing on per-user or scenario-specific differences. This choice reflects common real-world deployments, which must accommodate a wide range of behaviors and environments. Our approach involves five main stages:
  1. 1.
    We acquire large-scale sensor readings from publicly available datasets that capture diverse user activities and device usage patterns. These datasets encompass different motion, environmental, and orientation sensors. Detailed descriptions of each dataset are provided at the end of this section.
  2. 2.
    We merge sensor readings into a single, global distribution for each sensor modality in each dataset. For sensors that are inherently discrete or quantised (e.g., integer output ranges), we simply count occurrences. For sensors that produce (quasi-)continuous values, we rely on quantisation using Freedman–Diaconis binning to partition the output space and approximate an empirical probability mass function.
  3. 3.
    From these global distributions, we compute max, Shannon and collusion entropies to measure the best- and average-case uncertainties of sensor outputs, along with the min-entropy to characterise the worst-case unpredictability.
  4. 4.
    Many real-world proposals combine multiple sensor streams to purportedly increase security. To assess the impact on worst-case unpredictability, we use Chow-Liu trees to approximate the joint distributions of different sensor modalities. This allows us to estimate higher-dimensional entropies without incurring prohibitive computational costs. We discuss this in §4.3.
  5. 5.
    Finally, we interpret the resulting entropy measures, focusing on whether sensor outputs remain sufficiently unpredictable against an informed adversary. We compare single-sensor versus multi-sensor scenarios to verify if combining modalities truly alleviates biases or simply adds redundant data susceptible to similar predictability concerns.
Throughout this process, we remain mindful of well-documented constraints with NIST SP 800-90B, SP 800-22 [33], and AIS 20/31 [26] in analysing multi-sensor data streams [35, 17]. Such frameworks were not designed to analyse the joint entropy of complex, multivariate data sources, which is the aim of this work. (For example, NIST SP 800-90B focusses on assessing univariate entropy sources with reduced, i.e. 8-bit, output sizes [35, 33]). To begin with, we sought publicly available sensor datasets suitable for analysing motion and environmental data at scale. Our search involved broad queries across IEEE DataPort, Google Scholar, Google Dataset Search, and GitHub. Several ostensibly “open” datasets either were no longer downloadable or imposed restrictive licensing terms [36, 37, 38]. Ultimately, we narrowed our scope to four datasets that offer diverse usage contexts, consistent sampling rates, and documented sensor modalities:
  • 1.
    UCI-HAR [39]: A widely referenced dataset for human activity recognition, comprising smartphone sensor recordings from multiple subjects performing daily activities. Data includes triaxial accelerometer and gyroscope signals.
  • 2.
    University of Sussex–Huawei Locomotion (SHL) [40, 41]: sampled at 100 Hz from an Huawei Mate 9 smartphone. The publicly available SHL Preview dataset is used, comprising three recording-days per user (59 hours of data in total). To scope this study, we use the dataset from the handheld mobile phone as a good fit with related work.
  • 3.
    Relay [5]: Contains sensor measurements for approximately 1,500 NFC-based contactless transactions, each recorded at 100 Hz across several physical locations (e.g., cafés). The dataset contains accelerometer, gyroscope, and environmental readings from devices involved in emulated payment scenarios.
  • 4.
    PerilZIS [42]: Collected at 10 Hz from a Texas Instruments SensorTag, a Samsung Galaxy S6, and a Samsung Galaxy Gear, this dataset spans multiple zero-interaction security use cases in an office environment.
These four datasets provide a variety of sensor types, user activities, and sampling rates, allowing us to explore how intrinsic biases and correlations manifest across different scenarios. Next, we detail how we preprocess and aggregate this data to form global distributions for our entropy analyses.
4 Entropy Analysis
In this section, we analyse the intrinsic entropy of sensor data under the threat model described in §3. We begin by discussing the challenges in quantising naturally continuous sensor values for discrete-entropy calculations, then present our findings for our single- and multi-sensor analyses.
4.1 Pre-processing
A crucial, yet underexplored, issue in prior work (e.g. [15, 17, 16, 31]) is how to convert inherently continuous sensor outputs into suitable discrete values for entropy estimation. For example, Shannon and min-entropy, as defined in Eqs. 3 and 5, rely on discrete random variables. Physical quantities such as linear acceleration or angular velocity are continuous in nature, even though modern sensors employ internal analog-to-digital conversion with a finite resolution. Yet, a sensor’s advertised resolution (e.g. 12 bits for the widely used Bosch BMA mobile accelerometer [43]) does not imply uniform coverage across its range. Everyday usage introduces biases and clustering, resulting in some measurements occurring far more frequently than others. For instance, UCI-HAR data shows accelerometer readings concentrated in certain areas, and approximately 60% of gyroscope readings hover near zero (see Figure 1). Such skew and bias radically diminishes entropy compared to uniformly distributed values.
(a) Acc. $x$ axis.
(b) Acc. $y$ axis.
(c) Acc. $z$ axis.
(d) Gyro. $x$ axis.
(e) Gyro. $y$ axis.
(f) Gyro. $z$ axis.
(g) Accelerometer.
(h) Gyroscope.
(i) Light.
(j) Linear Acceleration.
(k) Magnetometer.
(l) Rotation Vector.
Figure 1: Global sensor data CDFs – UCI-HAR (a–f) and Relay (g–l) datasets.
[…]
$h=2\cdot\frac{IQR(x)}{n^{1/3}}$ (8)
4.2 Single Sensors
Table 1: Single-sensor entropy values (in bits) for each dataset. Grey cells denote unavailable data for that dataset and modality.
Dataset
UCI-HAR SHL Relay* PerilZIS
Sensor $H_{0}$ $H_{1}$ $H_{2}$ $H_{\infty}$ $H_{0}$ $H_{1}$ $H_{2}$ $H_{\infty}$ $H_{0}$ $H_{1}$ $H_{2}$ $H_{\infty}$ $H_{0}$ $H_{1}$ $H_{2}$ $H_{\infty}$
Acc.X 8.488 7.080 5.876 3.729 11.557 8.732 7.487 4.543 13.012 9.292 6.359 3.626
Acc.Y 8.243 7.231 6.847 5.694 11.425 8.928 7.717 4.500 9.549 5.873 4.483 2.889
Acc.Z 8.455 7.397 7.069 6.020 10.428 7.627 6.366 3.785 9.817 6.671 5.403 4.002
Acc.Mag 8.895 6.284 4.819 3.489 14.583 10.136 8.710 6.435 10.145 6.843 5.808 4.538 13.328 8.273 7.115 4.526
Gyro.X 8.683 5.430 3.504 1.929 15.024 10.532 8.107 4.993 14.528 7.231 4.454 2.805
Gyro.Y 8.439 5.023 3.461 2.300 15.085 10.283 7.601 4.827 14.078 6.715 4.039 2.529
Gyro.Z 8.714 5.675 3.948 2.363 15.281 10.070 5.708 3.083 13.961 6.463 3.836 2.434
Gyro.Mag 8.414 5.759 4.130 2.537 12.123 7.816 5.728 3.699 7.954 4.751 3.442 2.083 14.166 5.565 1.932 0.969
Mag.X 12.845 8.840 8.386 6.374 10.767 7.639 6.816 4.883
Mag.Y 12.263 8.737 8.314 6.223 10.179 7.622 6.605 4.405
Mag.Z 12.516 8.586 8.217 6.228 10.129 7.507 6.726 4.448
Mag.Mag 13.558 9.436 8.771 7.148 7.972 6.147 5.617 4.254 10.293 7.329 6.489 4.454
Rot. Vec. 8.725 7.721 5.970 3.220 5.000 3.307 1.965 1.021
Grav.X 9.014 8.482 7.266 4.299
Grav.Y 9.338 8.770 7.602 4.453
Grav.Z 8.180 7.193 5.418 3.036
Grav.Mag 14.373 7.988 7.227 6.242 7.794 6.325 5.864 4.532
LinAcc.X 15.260 10.077 7.621 5.116
LinAcc.Y 14.859 10.116 7.639 5.224
LinAcc.Z 14.377 9.951 7.605 4.543
LinAcc.Mag 12.777 7.968 5.752 3.420 9.175 6.385 5.424 4.222
Light 7.200 5.331 4.507 3.206 12.152 7.940 7.137 4.552
Humidity 7.943 7.048 6.774 5.546
Temp. 7.295 4.753 2.611 1.332 8.484 7.416 6.941 5.449
Pressure 9.461 8.170 7.723 6.237 8.044 7.006 6.370 5.073
Mean 8.541 6.235 4.957 3.508 13.188 9.266 7.178 4.483 7.891 5.584 4.661 3.408 11.277 7.224 5.717 3.912
S.D. 0.207 0.904 1.461 1.574 1.993 1.148 1.115 1.018 1.612 1.227 1.474 1.379 2.312 0.900 1.526 1.266
Given the biases discussed above, it is inevitable that some sensor readings will exhibit relatively high predictability. To quantify this, we calculate individual-sensor entropies across multiple datasets. The results are given in Table 1. For multi-dimensional modalities (e.g. triaxial accelerometer or gyroscope), these are split into separate axes following Voris et al. [15]. We note that, in the Relay dataset, the data for individual $x$, $y$ and $z$ components are not given for the accelerometer, gyroscope, and magnetometer sensors. Rather, the authors have already preprocessed triaxial data into its vector magnitudes, i.e. $\textbf{v}=\sqrt{x^{2}+y^{2}+z^{2}}$. We give this as “X.Mag” for a given sensor X. For completeness, we compute the magnitude ourselves for other datasets, where applicable, and report the entropy values for this new synthetic modality.
Several clear patterns emerge from Table 1. Some sensors, such as certain accelerometer axes in SHL or PerilZIS, exhibit moderate min-entropies of 4–6 bits. Other sensors, particularly gyroscope axes (see UCI-HAR) show values below 3 bits, indicating high predictability in their most frequent readings. Shannon entropy values ($H_{1}$) can be fairly high (up to 10 bits in some cases), whereas min-entropy ($H_{\infty}$) is often much lower. This gap reflects distributions where a few outcomes dominate, thereby driving worst-case unpredictability down even if the average-case picture is more favorable. Overall, the results confirm that data from individual sensors do not provide sufficient min-entropy for robust security on their own. In the next section, we examine whether combining multiple modalities can meaningfully increase this worst-case unpredictability or whether correlated biases persist across different sensor streams.
4.3 Multi-modal Sensors
Several sensor-based security proposals [13, 12, 4, 2, 3] assert that combining multiple sensor modalities can bolster security, based on the intuition that an adversary must accurately predict several data streams, rather than just one. This section will examine that claim.
(a) UCI-HAR
(b) PerilZIS
(c) Relay
(d) SHL
Figure 2: Sensor correlation matrices for each dataset.
A naïve approach might add Shannon entropies from individual sensors, benefitting from the relation $H(X_{1},\ldots,X_{n})=\sum_{i=1}^{n}H(X_{i})$. However, this requires that $X_{i}$ are statistically independent. In reality, mobile sensors often exhibit strong dependencies. For instance, the rotation vector, gravity, and linear acceleration sensors are frequently derived in software from the accelerometer and gyroscope on consumer devices [44]. As a result, these modalities cannot be treated as independent random variables. Figure 2 illustrates how multiple sensors in each dataset correlate: some pairs are nearly perfectly aligned (correlation close to ±$\pm 1$), which drastically reduces their combined unpredictability. High correlations invalidate the simplistic additive model of entropy. Even if multiple modalities individually appear to have moderate unpredictability, overlapping probability distributions may limit the overall joint entropy. In the next subsections, we discuss why straightforward joint-entropy calculations are computationally intractable at scale, before describing how Chow–Liu trees enable a practical approximation of higher-dimensional entropy.
4.3.1 Complexity Challenges
Computing the exact joint probability distribution and joint entropy of multiple sensors can quickly become prohibitively expensive. Let each of the $n$ sensor modalities be discretised into $b_{i}$ bins. Then, the joint distribution has $\prod_{i=1}^{n}b_{i}$ distinct states, an enormous state space once $n$ and $b_{i}$ grow. Applying Freedman–Diaconis binning rules typically results in thousands of bins per modality, causing the number of joint bins to explode combinatorially.
Moreover, even before enumerating states, selecting which sensors to combine can itself involve $2^{n}-(n+1)$ subsets, skipping single-sensor subsets and the empty set. Preliminary experiments confirmed joint entropies could be computed directly for $n\leq 3$ modalities with a maximum of 1250 bins and fewer than 150K total samples from the Relay dataset. Reducing bin sizes can help, but this risks oversimplifying the distribution and artificially deflating entropy estimates. Further experiments confirmed that limiting the bin numbers reduced our single-sensor entropy estimates by approximately 2–3 bits on average compared to those reported in Table 1. We therefore sought an alternative strategy that balances accuracy with tractable computation.
4.3.2 Chow-Liu Approximation
To handle these scaling issues, we adopt Chow–Liu trees [45], which approximate high-dimensional joint distributions using a maximum-weight spanning tree, $\pi$, over the different sensor modalities. Each edge is weighted by the mutual information of the connected variables, ensuring the tree structure captures the dominant pairwise dependencies. This approach minimises the Kullback–Leibler divergence (Def. 2.3) between the true multivariate distribution and the resulting tree-based approximation as follows:
$p_{\pi}(x_{1},\dots,x_{n})=p(x_{r})\prod_{i\neq r}p(x_{i}\;|\;x_{\pi(x)})$ (9)
Where $\pi(i)$ denotes the parent of $X_{i}$ in the tree, and $r$ is the tree’s root node. Chow–Liu trees are acyclic, singly connected structures: each node has at most one parent where one can traverse the tree to accumulate probabilities between pairwise dependencies. This significantly reduces computation time compared to naïve enumeration of the full joint measurement space. The use of Chow-Liu trees was proposed by Buller and Kaufer [35] for estimating the entropy of multivariate data sources where the range of possible values is high. In our Python implementation, we use the pgmpy [46] library’s TreeSearch module. Practically, for each sensor subset, we:
  1. 1.
    Discretise each sensor’s readings via Freedman–Diaconis binning.
  2. 2.
    Build a Chow–Liu tree from the mutual information of each sensor pair, selecting edges to form a spanning tree.
  3. 3.
    Traverse the resulting tree to estimate max ($H_{0}$), Shannon ($H_{1}$), collision ($H_{2}$), and min-entropy ($H_{\infty}$) without enumerating the full exponential state space.
Our framework evaluates the joint entropy over all sensor combinations. The powerset of the sensor set is generated and processed in parallel using Python’s multiprocessing module. Processing all datasets took approximately 22 hours on our workstation with an Intel i7-6700K (8M cache, 4.20 GHz) and 32 GB RAM on Ubuntu 24.04.
4.3.3 Results
Tables 25 report the top 10 performing multi-sensor combinations ranked by min-entropy for each dataset. As expected, combining all sensors yields the highest $H_{0}$ (max-entropy) and often increases Shannon and collision entropy. However, min-entropy ($H_{\infty}$) remains stubbornly low. For instance, the complete set of sensors in SHL surpasses 80 bits of $H_{1}$ (Shannon) but saturates at only 21 bits of $H_{\infty}$. Interestingly, we find that omitting certain correlated sensors sometimes does not reduce min-entropy at all. For the Relay dataset in Table 3, the combination (Acc., Gyro., Light, Lin. Acc., Mag., Rot. Vec.) achieves $H_{\infty}=7.859$ bits, only slightly below the full set’s $8.092$ bits. Parallel findings arise in the PerilZIS and SHL datasets, where omitting a small number of sensors from the “All sensors” set has negligible impact on $H_{\infty}$. This pattern appears across datasets: additional modalities may raise $H_{0}$ and $H_{1}$ but barely move $H_{\infty}$.
(a) UCI-HAR
(b) PerilZIS
(c) Relay
(d) SHL
Figure 3: Entropy values versus the number of joint sensor modalities.
[…]
The results suggest that many sensors contribute redundant information when combined, exposing a fundamental limitation of multi-modal data in reality. Combining signals increases the apparent capacity for unpredictability, but correlations between sensors means that the min-entropy from a large sensor ensemble is not be substantially higher than a reduced subset thereof. We note that standards implore the use of min-entropy as a safer, worst-case metric [26, 33]. It measures how close the distribution is to collapsing around the single most probable outcome that an adversary will target first. To the best of our knowledge, the entropy collapse from highly correlated sensor modalities has not been elicited in existing work. We provide the full sensor combination results in our open-source repository.2
Table 2: Top 10 best-performing sensor combinations (UCI-HAR; in bits).
Modality $H_{0}$ $H_{1}$ $H_{2}$ $H_{\infty}$
All sensors 48.061 25.089 17.116 11.008
(Acc.{x,y,z}, Gyro.{y,z}) 39.403 22.835 15.999 10.113
(Acc.{x,y,z}, Gyro.{x,y}) 39.886 21.456 15.332 9.900
(Acc.{x,y,z}, Gyro.{x,z}) 40.222 21.470 15.040 9.411
(Acc.{y,z}, Gyro.{x,y,z}) 40.228 21.306 14.698 9.274
(Acc.{x,z}, Gyro.{x,y,z}) 40.293 21.260 14.575 9.154
(Acc.{x,y,z}, Gyro.y) 31.228 18.804 13.893 9.065
(Acc.{x,y}, Gyro.{x,y,z}) 40.273 21.155 14.208 8.721
(Acc.{x,y,z}, Gyro.z) 31.564 18.775 13.833 8.682
(Acc.{y,z}, Gyro.{y,z}) 31.570 18.583 13.400 8.386
Table 3: Top 10 best-performing sensor combinations (Relay; in bits).
Modality $H_{0}$ $H_{1}$ $H_{2}$ $H_{\infty}$
All sensors 45.794 25.036 15.725 8.092
(Acc., Gyro., Light, Lin. Acc., Mag., Rot. Vec.) 44.795 24.963 15.334 7.859
(Acc., Grav., Light, Lin. Acc., Mag., Rot. Vec.) 38.385 21.590 14.783 7.702
(Acc., Grav., Gyro., Light, Mag., Rot. Vec.) 37.026 21.001 14.553 7.702
(Grav., Gyro., Light, Lin. Acc., Mag., Rot. Vec.) 36.092 20.794 14.450 7.673
(Acc., Light, Lin. Acc., Mag., Rot. Vec.) 37.385 21.517 14.428 7.469
(Acc., Gyro., Light, Mag., Rot. Vec.) 36.026 20.924 14.270 7.465
(Gyro., Light, Lin. Acc., Mag., Rot. Vec.) 35.092 20.720 14.147 7.439
(Acc., Grav., Gyro., Light, Lin. Acc., Mag.) 41.094 22.794 14.437 7.338
(Acc., Grav., Light, Mag., Rot. Vec.) 29.617 17.555 13.244 7.312
Table 4: Top 10 best-performing sensor combinations (PerilZIS; in bits).
Modalities $H_{0}$ $H_{1}$ $H_{2}$ $H_{\infty}$
All sensors 83.662 36.998 33.682 23.926
(Acc.{x,y,z}, Light, Temp., Pres., Mag.{x,y,z}, Gyro.{x,y,z}) 76.907 34.258 31.510 23.835
(Acc.{y,z}, Light, Temp., Pres., Mag.{x,y,z}, Gyro.{x,y,z}) 72.737 34.246 31.509 23.835
(Acc.{x,z}, Light, Temp., Pres., Mag.{x,y,z}, Gyro.{x,y,z}) 72.820 34.243 31.508 23.829
(Acc.{x,y}, Light, Temp., Pres., Mag.{x,y,z}, Gyro.{x,y,z}) 72.737 34.229 31.501 22.945
(Acc.{x,y,z}, Light, Temp., Mag.{x,y,z}, Gyro.{x,y,z}) 71.384 32.586 30.190 22.945
(Acc.y, Light, Temp., Press., Mag.{x,y,z}, Gyro.{x,y,z}) 68.568 34.218 31.500 22.945
(Acc.x, Light, Temp., Pres., Mag.{x,y,z}, Gyro.{x,y,z}) 68.650 34.215 31.499 22.945
(Acc.{y,z}, Light, Temp., Mag.{x,y,z}, Gyro.{x,y,z}) 67.214 32.575 30.189 22.945
(Acc.{x,y,z}, Light, Temp., Pres., Mag.{x,z}, Gyro.{x,y,z}, Hum.) 76.618 33.816 31.035 21.915
Table 5: Top 10 best-performing sensor combinations (SHL; in bits).
Modality $H_{0}$ $H_{1}$ $H_{2}$ $H_{\infty}$
All sensors 158.601 82.301 39.320 21.289
(Acc.{x,y,z}, Gyro.{x,y,z}, Mag.{x,y,z}, Ori.{w,x,y,z}, Grav.{x,y,z}, LinAcc.{y,z}, Pres., Alt., Temp.) 148.995 78.624 39.276 21.289
(Acc.{x,y,z}, Gyro.{x,y,z}, Mag.{x,y,z}, Ori.{w,x,y,z}, Grav.{x,y,z}, LinAcc.{x,z}, Pres., Alt., Temp.) 148.759 78.477 39.272 21.289
(Acc.{x,y,z}, Gyro.{x,z}, Mag.{x,y,z}, Ori.{w,x,y,z}, Grav.{x,y,z}, LinAcc.{x,y,z}, Pres., Alt., Temp.) 148.587 78.380 39.249 21.289
(Acc.{x,y,z}, Gyro.{x,y,z}, Mag.{x,y,z}, Ori.{w,x,y,z}, Grav.{x,y,z}, LinAcc.{x,y}, Pres., Alt., Temp.) 148.952 78.135 39.235 21.289
(Acc.{x,y,z}, Gyro.{x,y,z}, Mag.{x,y,z}, Ori.{w,x,y,z}, Grav.{x,y,z}, LinAcc.{z}, Pres., Alt., Temp.) 139.153 74.800 39.175 21.289
(Acc.{x,y,z}, Gyro.{x,z}, Mag.{x,y,z}, Ori.{w,x,y,z}, Grav.{x,y,z}, LinAcc.{y,z}, Pres., Alt., Temp.) 138.981 74.703 39.128 21.289
(Acc.{x,y,z}, Gyro.{x,z}, Mag.{x,y,z}, Ori.{w,x,y,z}, Grav.{x,y,z}, LinAcc.{x,z}, Pres., Alt., Temp.) 138.745 74.556 39.117 21.289
(Acc.{x,y,z}, Gyro.{x,y,z}, Mag.{x,y,z}, Ori.{w,x,y,z}, Grav.{x,y,z}, LinAcc.{y}, Pres., Alt., Temp.) 139.346 74.458 39.100 21.289
(Acc.{x,y,z}, Gyro.{x,y,z}, Mag.{x,y,z}, Ori.{w,x,y,z}, Grav.{x,y,z}, LinAcc.{x}, Pres., Alt., Temp.) 139.109 74.311 39.087 21.289
5 Evaluation
This section critically evaluates our sensor entropy findings using existing literature and relevant standards. We then discuss potential mitigations such as randomness extractors, outline inherent limitations of our study and sensor-based approaches, and propose avenues for future enhancements.
5.1 Discussion
Prior studies have claimed or implied that mobile sensors are suitable data sources for security-critical applications. Much work has relied on model evaluation metrics as a proxy for evaluating such claims [13, 14, 4, 2], with a smaller subset using more established entropy metrics [25, 24]. However, analyses have demonstrated that individual sensors confer very little entropy, especially under worst-case assumptions [17, 16, 15, 7]. Our results, derived from an analysis of multiple datasets and modalities (see Tables 25), corroborate and challenge these claims.
Single sensors yield low entropy. Building on prior work, we find that min-entropy often remains significantly below the Shannon estimates (e.g., some single-sensor readings yield only 1–3 bits). Our min-entropy estimates often remain significantly lower than their Shannon counterparts (e.g. some single-sensor readings yield only 1–3 bits of min-entropy). This exposes a major gap between average- and worst-case unpredictability across 25 widely used sensors, which is a critical distinction that existing schemes and evaluations do not fully capture.
Multiple sensors improve worst-case entropy, but not dramatically. While combining more modalities typically raises the upper bound (e.g. Hartley or Shannon entropy), our experiments reveal that min-entropy gains are far smaller than one might hope. In some cases, $H_{\infty}$ plateaus, indicating that a few highly probable outcomes still dominate the distribution. Our findings place stricter bounds on the benefits and risks of mobile-sensor-based approaches. We estimate that even the most complicated multi-modal combination provide relatively little worst-case entropy. It is important to note that our analysis does not even incorporate adversarial perturbations, meaning real-world attacks could degrade unpredictability even further. Hence, while mobile sensors can augment other authentication or key-generation processes, they rarely suffice as a standalone source. They might be useful for tasks where strong assurances are not required, e.g. simple proximity checks; however, relying on sensor data as robust entropy sources is fraught with security risks.
What can be done? One might hope that cryptographic extractors (e.g. Von Neumann extractors or more advanced schemes [47, 48, 49]) could make low-entropy sensors suitable for usage. Extractors are designed to reduce bias in a noisy or skewed source; however, they cannot increase the total amount of unpredictability beyond the source’s intrinsic min-entropy. If combined sensor data provides, say, 24 bits of min-entropy, then post-processing can at best produce a short, unbiased bitstring reflecting those 24 bits, and no more. This means that an extractor can improve the quality of the randomness (i.e., make it more uniform) but not increase the quantity (i.e., its total brute-force resistance). Extraction enhances the quality of the original source if it has enough min-entropy, but it cannot elevate a source that fundamentally lacks it.
Another avenue is to introduce additional unpredictability through user interaction or deliberate environmental perturbation. For example, requiring the user to perform a random shaking gesture during a pairing protocol could inject extra entropy into the sensor readings. Prior studies have found that deliberate device movements can, indeed, increase the usable entropy of motion sensors, albeit only by a modest amount (on the order of a few bits) [15, 17]. While this boost is non-negligible, it remains far below the dozens or hundreds of bits typically desired for security-critical applications.
One possibility is to incorporate sensor-based randomness into an entropy pool that seeds a cryptographically secure pseudo-random number generator (CSPRNG). By doing so, a hybrid design can mitigate bias in the raw sensor data and produce outputs that pass statistical randomness tests (see the designs by Suciu et al. [27] and Wallace et al. [29]). However, this approach implicitly contradicts the rationale for many existing sensor-based applications, which leverage similarity or correlation in sensor readings. For example, in matching similar motion traces or environmental conditions of two neighbouring devices [42, 13, 2, 3, 4, 1]. The desired property here is not pure unpredictability but rather shared information between the sensor signals collected by one or more devices. Using CSPRNG-derived signals would make it difficult to distinguish whether signals originated from co-located devices using similarity-based metrics. This exposes a fundamental security paradox in existing work. We have shown that sensors have insufficient entropy, even when using multiple ones simultaneously. Yet, methods to increase that entropy would compromise the usefulness of schemes that rely on the similarity of one or more sensor signals.
We see no straightforward remedy for the entropy shortfall of mobile sensors. Extractors, user-assisted randomness, and entropy pools address diHoweverfferent problems: reducing bias, injecting small amounts of fresh entropy, or improving output distribution. Nevertheless, they do not compensate for fundamentally weak signals, which we have shown to be the case across a range of modalities. Sensor data and its derivatives may still be appropriate for low-security purposes. However, relying on it alone for security applications is highly inadvisable due to the serious risk of adversaries who exploit statistical biases to reduce the effective search space.
5.2 Limitations
Despite our detailed analysis, this work has some notable limitations. Firstly, while we analyse four large datasets, they do not fully capture contexts all possible contexts. Some work has suggested that performing dedicated movements (e.g. gestures) can increase the amount of usable entropy from motion sensors by 5–6 bits [17, 15]. Our datasets do not cover such dedicated movements; it is possible that the reported results are an underestimation of entropy for sensors which are deliberately perturbed as an entropy-generating action. Secondly, our choice of Freedman-Diaconis binning and Chow-Liu trees is a pragmatic compromise. Smaller bin sizes tend to over-simplify the distribution and underestimate entropy, whereas larger bins can lead to computational blowup. As such, although our approach outperforms naïve joint entropy estimation, it is still an approximation. Thirdly, we largely focus on a data-centric view of entropy and do not examine other threats, such as sensor spoofing [3], hardware attacks [34] and cross-device correlation attacks (see [50] against wireless body area networks). It is possible that these threats may reduce the entropy of sensors even further.
6 Conclusion
This paper provides a comprehensive analysis of sensor-derived entropy across multiple datasets and modalities. Our results expose a tension between the perceived and actual strength of sensor data for security applications. Even in the best-performing sensor combinations, seemingly suitable results using one metric collapse to insecure levels when using standard worst-case metrics. Notably, modalities that yield ‘good’ max- or Shannon entropies, representing the best- and average-case unpredictability, have insecure worst-case min-entropies. Consequently, sensor modalities that may appear robust have biases that may enable adversaries to predict the most probable values with minimal effort.
Our findings also challenge the prevailing orthodoxy that model evaluation metrics (e.g. accuracy or EER) suffice to demonstrate the inherent randomness of sensor signals. The vulnerability of mobile sensors to biased distributions significantly undermine their effectiveness as reliable entropy sources. We also cast doubt on the use of sensors with respect to their non-stationary and lack of reproducibility. These issues collectively contradict the criteria articulated in frameworks such as NIST SP 800-90B, which emphasise noise-source stationarity and protection from external influence. The effectiveness of countermeasures remains an open research challenge. Our hope is that the methodologies and insights presented here will encourage the security community to adopt more rigorous evaluation strategies for sensor-based techniques, paving the way for safer and more robust designs in mobile device security.
In future work, we consider that a dynamic analysis is important to assess the stationarity issues with sensor data, where entropy varies according to user behavior or environment. Moreover, a user study could yield empirical bounds on how finely humans can intentionally manipulate motion sensors, revealing more realistic limits to sensor-based randomness in real-world scenarios. For example, our binning decision was a statistical one, rather one that reflects human usage; it is possible that real-world usage may reduce the resolution of useful sensor data, thus reducing entropy. Overall, while sensor-based data can reach limited levels of entropy under favourable conditions, the road to making such sources systematically secure and robust is long. The key takeaway is that substantial work is needed before sensors can be considered appropriate for security-critical applications.
Acknowledgments
Carlton Shepherd has received funding from the UK EPSRC ‘Chameleon’ project (EP/Y030168/1). The authors would like to thank Darren Hurley-Smith for insightful conversations on the topic.