### v8

# Entropy Collapse in Mobile Sensors: The Hidden Risks of Sensor-Based Security

Abstract
Mobile sensor data has been proposed for security-critical applications such as device pairing, proximity detection, and continuous authentication. However, the foundational premise that these signals provide sufficient entropy remains under-explored. In this work, we systematically analyse the entropy of mobile sensor data using four datasets from multiple application contexts (UCI-HAR, SHL, Relay, and PerilZIS). Using direct computation and estimation, we report entropy values—max, Shannon, collision, and min-entropy—for an exhaustive range of sensor combinations. We demonstrate that the entropy of mobile sensors remains far below what is considered secure by modern standards for security applications, such as for authentication and key generation systems, even when many sensors are combined. In particular, we observe an alarming divergence between average-case Shannon entropy and worst-case min-entropy. Single-sensor mean min-entropy varies between 3.408–4.483 bits despite Shannon entropy being several multiples higher. We further show that correlations between sensor modalities contribute to a $\approx$75% reduction between Shannon and min-entropy. This brings joint min-entropy well below 10 bits in many cases and, in the best case, yielding only $\approx$24 bits when combining 20+ sensor modalities. Our results reveal that adversaries may feasibly predict sensor signals through an exhaustive exploration of the measurement space. Our work also calls into question the widely held assumption that adding more sensors inherently yields higher security. Ultimately, we strongly urge caution when relying on mobile sensor data for security applications.
keywords:
Entropy , Sensors , Mobile security
1 Introduction
Modern mobile devices expose rich motion and ambient signals from accelerometers, gyroscopes, magnetometers, and more. Beyond activity recognition and context inference, e.g. [1, 2, 3, 4], a long line of work proposes using these signals for security-critical tasks. This includes cryptographic key generation, zero-interaction device pairing, and continuous authentication [5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20]. Traditional key generation protocols [8, 18] have historically relied on shared motion patterns in order to establish secure communication between co-located devices. Meanwhile, other proposals have incorporated environmental phenomena, such as characteristics of magnetic fields and thermal fluctuations, to mitigate relay attacks [9, 10, 21, 5, 11] and to reduce the frequency of user authentication prompts [17, 16, 12, 13, 14].
Despite these advances, a crucial assumption remains under-examined: that sensor data provides sufficient entropy to thwart adversarial inference. Many existing studies rely on heuristic assessments or model classsification metrics, e.g. [21, 7, 6, 9, 22, 19, 20], that do not address two basic questions: (1) how much entropy do sensors actually provide, and (2) to what extent do multimodal sensor combinations improve entropy? For the latter, multimodal sensing has been advocated to counter sensor-specific weaknesses [9, 21, 22, 7], but the quantitative security benefits of combining multiple sensors has not been rigorously evaluated.
The security implications of low-entropy sensor data can be direct and severe. If such data is used as authentication evidence or to derive secret keys, an adversary may attempt to enumerate the entire space of possible measurement values in order to (re-)derive it. For such reasons, modern security standards provide important quantitative thresholds for what can be considered ‘secure’. For instance, NIST SP 800-57 [23] states that security levels below 80 bits are broken for long-term use, with 112 bits as the current minimum and 128 bits recommended for security keys. Standards that govern dedicated entropy sources are even stricter: NIST SP 800-90B [24] expects entropy sources used for cryptographic random number generators (RNGs) to approach full per-bit entropy, [^2] More specifically, $\approx$1 bit ($\epsilon\leq 2^{-32}$ per-bit) of min-entropy per output bit. while AIS 20/31 [25] requires an RNG’s internal state to possess over 240 and 250 bits of min-entropy and Shannon entropy respectively. As we will demonstrate, the min-entropy of even complex combinations of mobile sensors falls short of these standards.
In this paper, we analyse the entropy of 25 mobile sensors, which is substantially more than prior work—[26] (1 sensor), [27] (10), [17] (2), [28] (3)—across multiple public datasets. Our results reveal structural limits. Commodity sensors exhibit pronounced biases and dependencies, yielding on average 3.408–4.483 bits of min-entropy and 5.584–9.266 bits of Shannon entropy per sample across datasets. Multimodal fusion confers only modest gains: inter-sensor correlations and statistical bias both contribute to a divergence between average- and worst-case entropy, with min-entropy typically 40–75% below Shannon entropy for the same signals. In practical terms, schemes that treat these measurements as secrets expose a guess space small enough to be systematically enumerable. Our contributions are as follows:
  • 1.
    We introduce the first systematic approach to evaluating sensor entropy across such a comprehensive range of modalities and datasets using various entropy metrics (max, Shannon, collision, and min-entropy).
  • 2.
    We empirically demonstrate how inter-sensor correlations and biases erode entropy, casting doubt on the proposition that using multiple sensors adds substantially to security.
  • 3.
    We show how the collapse in worst-case entropy opens the door to attacks that exhaustively enumerate the space of possible measurements. This raises fundamental concerns with respect to security applications that rely on signals from single and multiple mobile sensors. Consequently, we do not recommend relying wholly on mobile sensors as entropy sources for security-critical applications.
The rest of this paper is organised as follows: §2 discusses sensor-based security mechanisms and established entropy metrics. §3 explains our experiment design for entropy estimation, including the threat model and dataset selection. §4 presents empirical results of our analyses and §5 discusses the implications for system design. We conclude in §6 with recommendations for further work. Our analysis work is released publicly to foster future research. [^3] https://github.com/cgshep/entropy-collapse-mobile-sensors
2 Background
This section discusses sensor-based applications and existing approaches, and formalises the entropy metrics underpinning our subsequent analyses.
2.1 Related Work
Mobile sensing has been widely explored for security tasks in three broad families: (1) proximity detection and relay-attack mitigation, where co-presence is inferred from shared motion or ambient context [29, 7, 11, 5, 21, 6, 9]; (2) device pairing, where two devices derive evidence of co-location to establish a secure association or shared secret [8, 30, 18, 20]; and (3) continuous authentication, which aims to verify users passively during normal interaction [16, 12, 13, 14, 31, 32, 33, 34]. Representative designs include tap-to-pair schemes using accelerometers [7], induced shared vibrations [6], and ambient-signal approaches that leverage temperature, humidity, magnetic field, or other environmental cues [9, 21]. These mechanisms commonly fuse accelerometer and gyroscope measurements to capture correlated motion as evidence of co-presence [8, 30].
Our key observation is that model-performance metrics are typically decoupled from the entropy of the underlying data sources upon which such models are based. Across the aforementioned application areas, effectiveness is usually inferred from discriminative performance. A common pipeline—documented across proximity detection, pairing, and continuous authentication—collects data from $N$ participants; derives time- or frequency-domain features (e.g. cross-correlation, spectral energy, and Hamming, Euclidean, and mean absolute distances); learns a decision rule (thresholds or supervised models such as SVMs, Random Forests, Naïve Bayes); and reports FPR/FNR, EER, ROC/AUC, accuracy, or F1. This template appears in numerous systems, including [7, 9, 21, 11, 16, 14, 13, 31, 32, 6]. For a comprehensive treatment of the model evaluation metrics used by existing mobile sensor-based security systems, the reader is referred to the surveys by Abuhamad et al. [19], Alzubaidi and Kalita [35], and Mayrhofer and Sigg [20]. While discriminative performance metrics are informative for usability and functionality, they do not quantify entropy of the underlying sources, which is the focus of our analysis.
Adjacent evidence also highlights practical limits under tight timing constraints. For NFC transactions, EMV specifies $\approx$500 ms for proximity checks. In that setting, Markantonakis et al. [21] (building on [5, 11]) report EERs of 0.179–0.246 across sensor combinations, concluding that commodity sensors are unsuitable for reliable proximity and relay detection without incurring usability or security issues. Taken together, the literature demonstrates functional feasibility but leaves open the core question of how unpredictable the underlying measurements really are, especially under multimodal fusion. This motivates our entropy-centric evaluation in the remainder of the paper.
Some proposals have attempted to assess the entropy of sensor signals within the context of a security mechanism, but this represents a small minority of the literature. Specifically, T2Pair by Li et al. [15] reports 32.3–38.5 bits of Shannon entropy, and a refinement by Wu et al. [36] reports 51–54 bits. Even in the best cases, these values are low relative to modern cryptographic standards; moreover, we are aware of no work using standard min-entropy as a worst-case measure of unpredictability [24, 25, 37] (see comparison tables of state-of-the-art schemes in [35, 18, 20]). We address this gap by analysing the entropy of a wide range of mobile sensors, both in isolation and in combination.
Concurrent work has attempted to assess the entropy bestowed by mobile sensors as a separate line of inquiry. Voris et al. [26] investigated accelerometers as true random number generators (TRNGs) on a WISP RFID tag and Nokia N97 phone. The authors find that min-entropy is proportional to the motion applied to the device, with stationary movement having the lowest min-entropy. Intrinsic noise from the sensor’s circuitry and seismic activity, and the sampling rate of its analog-to-digital converter (ADC), are found to influence entropy generation. Min-entropy values of 3.1–11.4 bits were measured depending on the movement of the accelerometer. Lv et al. [28] analysed three mobile sensors—a triaxial accelerometer, gyroscope, and magnetometer—on an undisclosed Xiaomi Redmi smartphone. Min-entropy values of 0.593–5.876 are reported, depending on the modality and the entropy estimation method. Krhovják et al. [17] examined image and audio data collected from mobile phone cameras and microphones respectively on a Nokia N73 and E-Ten X500 and M700 phones. The work reports Shannon entropies of 2.9 (microphone) and 2.408–5.376 (camera), and min-entropies of 0.5 (microphone) and 0.754–3.928 (camera). Hennebert et al. [27] analysed 10 sensors on two wireless sensor monitors: a TI eZ430-RF2500 and a Zolertia Z1. A single-sensor analysis is presented, yielding min-entropy values of 0–7.85; motion sensors, e.g. accelerometer and vibration sensors, produced the highest entropy.
Sensors have also been suggested as entropy sources in low-cost RNG designs for mobile devices. Suciu et al. [38] proposed combining GPS along with accelerometer, gyroscope and orientation sensors on an HTC Google Nexus One, passing NIST SP 800-22 tests [39], but without precise entropy estimates. Wallace et al. [40] explored a multi-sensor RNG design (accelerometer, gyroscope, microphone, WiFi, GPS, camera) with a non-standard entropy evaluation using 37 Android devices. Unlike dedicated TRNGs (e.g. ring oscillators, Johnson–Nyquist thermal noise, quantum sources [41]), mobile sensors depend heavily on user behaviour and environmental factors, introducing biases and correlations absent in controlled entropy sources. Existing sensor-based mechanisms largely overlook these dynamics, relying principally on heuristic or model evaluation metrics that do not account well for skewed distributions and other biases affecting predictability. In contrast, proposals of new entropy sources and authentication mechanisms commonly rely on worst-case metrics, such as min-entropy as stipulated in NIST SP 800-90B and AIS 20/31 [42, 43, 41, 24, 25]. Our study bridges the gap by systematically evaluating entropy across modalities and datasets.
Table 1: Comparative summary of prior sensor entropy analyses.
Reference Sensorsa Metricsb Results (bits) Context / Use-Case
Voris et al. [26] Acc. $H_{\infty}$ 3.1–11.4 True random number generation
Hennebert et al. [27] 10 (Various) $H_{\infty}$ 0–7.85 Harvesting entropy from physical mobile sensors
Lv et al. [28] Acc., Gyr., Mag. $H_{\infty}$ 0.593–5.876 Entropy source analysis on smartphone
Krhovják et al. [17] Cam., Mic. $H_{\{1,\infty\}}$ 0.5–5.376 Randomness sources in mobile devices
Li et al. [15] Mic., Light $H_{1}$ 32.3–38.5 Zero-interaction device pairing
Wu et al. [36] Mic., Light $H_{1}$ 51–54 Zero-interaction device pairing
This work 25 (Various) $H_{\{0,1,2,\infty\}}$ 8.1–23.9 ($H_{\infty}$) Systematic entropy analysis, including sensor combinatons, for security systems
  • a
    Acc.: Accelerometer; Gyr.: Gyroscope; Mag.: Magnetometer; Cam.: Camera; Mic.: Microphone.
  • b
    $H_{0}$: Max; $H_{1}$: Shannon; $H_{2}$: Collision; $H_{\infty}$: Min-entropy.
In general, while several studies have investigated the entropy of specific sensors, a direct comparison is often difficult due to varying methodologies and contexts. Table 1 provides a structured summary of key prior works, contrasting the sensors analysed, entropy metrics used, and the reported entropy ranges.
2.2 Definitions
We use the following definitions and notation throughout this work. Furthermore, unless stated otherwise, all logarithms are base 2 and entropies are reported in bits.
Definition 2.1 (Rényi Entropy).
Let $X$ be a discrete random variable taking values in a set $\mathcal{X}$ with probability mass function $p(x)$. The Rényi entropy of order $\alpha$ ($\alpha>0$, $\alpha\neq 1$) is defined as
$H_{\alpha}(X)\,=\,\frac{1}{1-\alpha}\,\log\!\Biggl{(}\sum_{x\in\mathcal{X}}p(x)^{\alpha}\Biggr{)}.$ (1)
We draw attention to four special cases of $\alpha$ that are widely used in the literature. First, the Hartley (Max) Entropy (Eq. 2) is the logarithm of the number of possible outcomes with non-zero probability and serves as an upper bound.
$H_{0}(X)\equiv\lim_{\alpha\to 0}H_{\alpha}(X)\,=\,\log\Bigl{|}\bigl{\{}\,x\in\mathcal{X}:p(x)>0\bigr{\}}\Bigr{|}.$ (2)
Second, Shannon entropy (Eq. 3) is the classical definition of entropy in information theory, corresponding to the limit of $H_{\alpha}$ as $\alpha\to 1$.
$H_{1}(X)\equiv\lim_{\alpha\to 1}H_{\alpha}(X)\,=\,-\sum_{x\in\mathcal{X}}p(x)\,\log p(x).$ (3)
Third, the collision entropy ($\alpha=2$) quantifies the probability of drawing the same value twice from $X$:
$H_{2}(X)\,=\,-\log\Bigl{(}\sum_{x\in\mathcal{X}}p(x)^{2}\Bigr{)}.$ (4)
We note that $H_{1}$ provides an average-case measure of uncertainty. It takes into account the entire distribution of outcomes; however, an adversary may only need to guess the most likely event to gain an advantage. Min-entropy (Eq. 5) is used as a conservative, worse-case metric, accounting for the least favorable distribution of outcomes; it is the value in the limit $\alpha\to\infty$:
$H_{\infty}(X)\equiv\lim_{\alpha\to\infty}H_{\alpha}(X)\,=\,-\log\Bigl{(}\max_{x\in\mathcal{X}}p(x)\Bigr{)}.$ (5)
$H_{\alpha}$ is a non-increasing function of $\alpha$, i.e. $H_{\infty}(X)\leq\dots\leq H_{1}(X)\leq H_{0}(X)$. $H_{\infty}(X)$ focuses on the single most likely outcome, providing a strictly tighter (and generally minimum) bound on uncertainty. We observe that min-entropy ensures that even the most skewed probability distributions still meet the required security guarantees; indeed, it is a recommended method for assessing entropy sources in modern standards [24, 25]. We also rely on joint entropy for assessing the entropy of multiple random variables.
Definition 2.2 (Joint Rényi Entropy).
Let $X_{1},X_{2},\dots,X_{n}$ be discrete random variables that jointly take values in $\mathcal{X}_{1}\times\mathcal{X}_{2}\times\cdots\times\mathcal{X}_{n}$, with joint probability mass function $p(x_{1},x_{2},\dots,x_{n})$. The Rényi entropy of order $\alpha$ ($\alpha>0,\alpha\neq 1$) for these $n$ variables is defined as the following, where the sum is taken over all $(x_{1},\dots,x_{n})$ in $\mathcal{X}_{1}\times\mathcal{X}_{2}\times\cdots\times\mathcal{X}_{n}$:
$H_{\alpha}(X_{1},X_{2},\dots,X_{n})\,=\,\frac{1}{1-\alpha}\,\log\Biggl{(}\sum_{(x_{1},x_{2},\dots,x_{n})}p(x_{1},x_{2},\dots,x_{n})^{\alpha}\Biggr{)}$ (6)
In the limit $\alpha\to 1$, then $H_{\alpha}(X_{1},X_{2},\dots,X_{n})$ converges to the classical joint Shannon entropy of these $n$ variables. To support the later discussion on Chow-Liu trees—our approach to joint entropy estimation—we define the Kullback-Leibler divergence.
Definition 2.3 (Kullback-Leibler (KL) Divergence).
Given a true probability distribution $P(x)$ of a random variable and an approximate or reference distribution $Q(x)$, the KL divergence is defined as follows:
$D_{KL}(P\;||\;Q)=\sum_{x\in\mathcal{X}}P(x)\log\Big{(}\frac{P(x)}{Q(x)}\Big{)}$ (7)
3 Experiment Design
3.1 Threat Model and Assumptions
We consider an adversary aiming to compromise sensor-based security schemes (e.g., key generation, proximity detection, and continuous authentication). The adversary may gather extensive statistical data about how smartphone sensors behave in everyday usage; for instance, from widely available open datasets. We assume the attacker focuses on predicting or guessing the sensor outputs by prioritising the most probable values first, exploiting any biases in the distribution of sensor measurements. As such, they may resort to exhaustive enumeration of the measurement space if the target source’s entropy is low enough. We exclude capabilities such as fault injection and other hardware attacks (see [44]) as well as side-channel attacks such as cross-device correlation attacks [45]. While those could further reduce the effective entropy space—say, by inducing errors in the output values of sensing hardware—we regard them as outside the scope of this work. This threat model represents an adversary who can capitalise on statistical biases in sensor data without directly comprising the device physically. Our goal is to evaluate whether sensor data distributions offer sufficient entropy to resist attacks that search the space of sensor measurement values informed by their statistical properties.
3.2 High-level Methodology
In this work, we focus on estimating population-level (“global”) sensor distributions by pooling all users within a dataset. This choice reflects typical deployments in which systems must serve previously unseen users (e.g. first-time proximity checks, payment defence mechanisms, or ad-hoc device pairing) and thus cannot assume per-user enrollment or stable user-specific priors. Our approach involves five main stages:
  1. 1.
    Data acquisition. We collect large-scale sensor readings from public datasets spanning diverse activities and device usage. Modalities include motion, environmental, and orientation sensors (dataset details appear at the end of this section).
  2. 2.
    Global distributions. We merge sensor readings into a single, global distribution for each sensor modality in each dataset. For sensors that are inherently discrete or quantised (e.g., integer output ranges), we simply count occurrences. For sensors that produce (quasi-)continuous values, we rely on quantisation using Freedman–Diaconis binning to partition the output space and approximate an empirical probability mass function.
  3. 3.
    Single-sensor entropy analysis. From these global distributions, we compute max, Shannon and collusion entropies to measure the best- and average-case uncertainties of sensor outputs, along with the min-entropy to characterise the worst-case unpredictability.
  4. 4.
    Multimodal entropy estimation. Many real-world proposals combine multiple sensor streams to purportedly increase security. To assess the impact on worst-case unpredictability, we use Chow-Liu trees to approximate the joint distributions of different sensor modalities. This allows us to estimate higher-dimensional entropies without incurring prohibitive computational costs (discused in §4.3).
  5. 5.
    Interpretation. Finally, we interpret the resulting entropy measures, focusing on whether sensor outputs remain sufficiently unpredictable against an informed adversary. We compare single-sensor versus multi-sensor scenarios to verify if combining modalities truly alleviates biases or simply adds redundant data susceptible to predictability concerns.
Throughout, we note the constraints of NIST SP 800-90B, SP 800-22 [24], and AIS 20/31 [25] when applied to multi-sensor streams [37, 28]. These frameworks target univariate sources rather than joint entropy of complex multivariate processes, which is the focus of our work. (For example, NIST SP 800-90B focusses on assessing entropy sources with reduced, i.e. 8-bit, outputs [37, 24]). To scope suitable data, we searched IEEE DataPort, Google Scholar, Google Dataset Search, and GitHub. Several ostensibly “open” datasets were unavailable for download or had restrictive licensing [46, 47, 48]. We selected four datasets offering diverse contexts, consistent sampling, and documented modalities:
  • 1.
    UCI-HAR [3]: A widely referenced dataset for human activity recognition, comprising smartphone sensor recordings from multiple subjects performing daily activities. Data includes triaxial accelerometer and gyroscope signals.
  • 2.
    University of Sussex–Huawei Locomotion (SHL) [1, 2]: Sampled at 100 Hz from an Huawei Mate 9 smartphone. TWe use the publicly available SHL Preview dataset (three recording days per user; 59 hours total) and, for scope alignment with prior work, the handheld-phone subset.
  • 3.
    Relay [5]: Approximately 1,500 NFC-based contactless transactions recorded at 100 Hz across several physical locations (e.g. cafés), including accelerometer, gyroscope, and environmental readings in emulated payment scenarios.
  • 4.
    PerilZIS [49]: Collected at 10 Hz from a TI SensorTag, Samsung Galaxy S6, and Samsung Galaxy Gear, spanning multiple zero-interaction security use cases in an office environment.
These datasets provide varied sensor types, user activities, and sampling rates, enabling us to observe how biases and inter-sensor correlations manifest across scenarios. Next, we describe preprocessing and aggregation to form the global distributions used in our entropy analyses.
4 Entropy Analysis
In this section, we analyse the entropy of sensor data on a single- and multi-modal basis. We begin by discussing the challenges in quantising sensor values for discrete-entropy calculations, then present our findings for our single- and multi-sensor analyses.
4.1 Pre-processing
A crucial, yet underexplored, issue in prior work (e.g. [26, 28, 27, 42]) is how to appropriately estimate entropy from quantised sensor outputs. Physical quantities such as linear acceleration or angular velocity are continuous in nature, even though modern sensors employ internal analog-to-digital conversion with a finite resolution. Yet, a sensor’s advertised resolution—for instance, 12 bits in the widely used Bosch BMA400 mobile accelerometer [50]—does not imply uniform histograms across its range of values. Everyday usage produces highly non-uniform biases and clustering, resulting in some measurements occurring more frequently than others. For example, UCI-HAR data shows accelerometer readings concentrated in certain areas, with approximately 60% of gyroscope readings concentrated around zero. As an illustration, Figure 1 shows the cumulative probability distributions of some selected modalities across our datasets. We note that such skew and bias radically diminishes entropy compared to uniformly distributed values.
(a) Acc. $x$ axis.
(b) Acc. $y$ axis.
(c) Acc. $z$ axis.
(d) Gyro. $x$ axis.
(e) Gyro. $y$ axis.
(f) Gyro. $z$ axis.
(g) Accelerometer.
(h) Gyroscope.
(i) Light.
(j) Linear Acceleration.
(k) Magnetometer.
(l) Rotation Vector.
Figure 1: Global sensor data CDFs – UCI-HAR (a–f) and Relay (g–l) datasets.
Another practical issue is the sparsity of sensor measurements at high resolutions. Treating every minute fluctuation e.g. $9.001\,\mathrm{m\,s^{-2}}$ vs. $9.002\,\mathrm{m\,s^{-2}}$) as distinct outcomes can inflate entropy estimates. In real-world applications, it is the ‘similarity’ between measurement signals that is considered useful in existing work; humans cannot reliably reproduce high-precision movements capable of utilising a sensor’s digital resolution at, say, a $0.01\mathrm{ms^{-2}}$ scale. We therefore aggregate adjacent codes/values into bins to obtain reliable probability mass functions (pmfs). The choice of bin width is non-trivial: classical rules (e.g. Sturges, Scott, Doane) encode different assumptions about the underlying distribution, vary in computational cost, and differ in robustness to outliers. We adopt the Freedman–Diaconis (FD) rule, a distribution-agnostic, outlier-robust choice that adapts to sample size and spread, [^4] Alternatively, a perceptual/actuation-driven scheme aligned with human reproducibility could be used; we defer this to future work. with width $h$ given in Eq. 8, where $\mathrm{IQR}(x)$ is the interquartile range and $n$ the number of samples. Later, in §4.3.4, we show that FD falls within the plateau region as part of a bin sensitivity analysis, yielding stable estimates across datasets while avoiding oversmoothing and sparsity.
$h=2\cdot\frac{IQR(x)}{n^{1/3}}$ (8)
4.2 Single Sensors
Table 2: Single-sensor entropy values (in bits) for each dataset. Grey cells denote unavailable data for that dataset and modality.
Dataset
UCI-HAR SHL Relay* PerilZIS
Sensor $H_{0}$ $H_{1}$ $_{2} HH_{\infty}$ $H_{0}$ $H_{1}$ $_{2} HH_{\infty}$ $H_{0}$ $H_{1}$ $_{2} HH_{\infty}$ $H_{0}$ $H_{1}$ $_{2} HH_{\infty}$
Acc.X 8.488 7.080 5.876 3.729 11.557 8.732 7.487 4.543 13.012 9.292 6.359 3.626
Acc.Y 8.243 7.231 6.847 5.694 11.425 8.928 7.717 4.500 9.549 5.873 4.483 2.889
Acc.Z 8.455 7.397 7.069 6.020 10.428 7.627 6.366 3.785 9.817 6.671 5.403 4.002
Acc.Mag 8.895 6.284 4.819 3.489 14.583 10.136 8.710 6.435 10.145 6.843 5.808 4.538 13.328 8.273 7.115 4.526
Gyro.X 8.683 5.430 3.504 1.929 15.024 10.532 8.107 4.993 14.528 7.231 4.454 2.805
Gyro.Y 8.439 5.023 3.461 2.300 15.085 10.283 7.601 4.827 14.078 6.715 4.039 2.529
Gyro.Z 8.714 5.675 3.948 2.363 15.281 10.070 5.708 3.083 13.961 6.463 3.836 2.434
Gyro.Mag 8.414 5.759 4.130 2.537 12.123 7.816 5.728 3.699 7.954 4.751 3.442 2.083 14.166 5.565 1.932 0.969
Mag.X 12.845 8.840 8.386 6.374 10.767 7.639 6.816 4.883
Mag.Y 12.263 8.737 8.314 6.223 10.179 7.622 6.605 4.405
Mag.Z 12.516 8.586 8.217 6.228 10.129 7.507 6.726 4.448
Mag.Mag 13.558 9.436 8.771 7.148 7.972 6.147 5.617 4.254 10.293 7.329 6.489 4.454
Rot. Vec. 8.725 7.721 5.970 3.220 5.000 3.307 1.965 1.021
Grav.X 9.014 8.482 7.266 4.299
Grav.Y 9.338 8.770 7.602 4.453
Grav.Z 8.180 7.193 5.418 3.036
Grav.Mag 14.373 7.988 7.227 6.242 7.794 6.325 5.864 4.532
LinAcc.X 15.260 10.077 7.621 5.116
LinAcc.Y 14.859 10.116 7.639 5.224
LinAcc.Z 14.377 9.951 7.605 4.543
LinAcc.Mag 12.777 7.968 5.752 3.420 9.175 6.385 5.424 4.222
Light 7.200 5.331 4.507 3.206 12.152 7.940 7.137 4.552
Humidity 7.943 7.048 6.774 5.546
Temp. 7.295 4.753 2.611 1.332 8.484 7.416 6.941 5.449
Pressure 9.461 8.170 7.723 6.237 8.044 7.006 6.370 5.073
Mean 8.541 6.235 4.957 3.508 13.188 9.266 7.178 4.483 7.891 5.584 4.661 3.408 11.277 7.224 5.717 3.912
S.D. 0.207 0.904 1.461 1.574 1.993 1.148 1.115 1.018 1.612 1.227 1.474 1.379 2.312 0.900 1.526 1.266
Given the biases discussed above, it is inevitable that some sensor readings will be highly predictable. To quantify this, we calculate individual-sensor entropies across datasets. The results are given in Table 2. For multi-dimensional modalities (e.g. triaxial accelerometer or gyroscope), these are split into separate axes following Voris et al. [26]. This approach allows for a fine-grained analysis of directional biases, as movement along one axis may be more constrained or predictable than others in typical usage scenarios. We note that, in the Relay dataset, the data for individual $x$, $y$ and $z$ components are not given for the accelerometer, gyroscope, and magnetometer sensors. Rather, the authors have already preprocessed triaxial data into its vector magnitudes, i.e. $\textbf{v}=\sqrt{x^{2}+y^{2}+z^{2}}$. This magnitude represents the overall intensity of motion, irrespective of direction. We give this as “X.Mag” for a given sensor X. For the sake of completeness and consistency, we compute the magnitude for other datasets, where applicable, and report the entropy values for this synthetic modality.
Several clear patterns emerge from Table 2. Some sensors, such as certain accelerometer axes in SHL or PerilZIS, exhibit moderate min-entropies of 4–6 bits. Other sensors, particularly gyroscope axes (see UCI-HAR) show values below 3 bits, indicating high predictability. Shannon entropy values ($H_{1}$) can be fairly high (up to 10 bits in some cases), whereas min-entropy ($H_{\infty}$) is often much lower. This gap reflects distributions where a few outcomes dominate, driving worst-case unpredictability down even if the average-case picture is more favorable. Overall, the results confirm that data from individual sensors do not provide sufficient min-entropy for robust security on their own. In the next section, we examine whether combining multiple modalities can meaningfully increase this worst-case unpredictability or whether correlated biases persist across different sensor streams.
4.3 Multi-modal Sensors
Several sensor-based security proposals [7, 22, 21, 9, 10] assert that combining multiple sensor modalities bolsters security, on the intuition that an adversary must accurately predict several data streams rather than one. Let us now explore that claim. Initially, a naïve approach might add Shannon entropies from individual sensors, benefitting from the relation $H(X_{1},\ldots,X_{n})=\sum_{i=1}^{n}H(X_{i})$. However, this requires that $X_{i}$ are statistically independent. In reality, mobile sensors often exhibit strong dependencies; for instance, the rotation vector, gravity, and linear acceleration sensors are frequently derived in software from the accelerometer and gyroscope on consumer devices [51]. As a result, these modalities cannot be treated as independent variables.
(a) UCI-HAR
(b) PerilZIS
(c) Relay
(d) SHL
Figure 2: Pairwise sensor correlation matrices for each dataset.
Figure 2 provides a quantitative view of inter-sensor redundancy through Pearson correlation matrices for each dataset. Strong correlations, with $|\rho|\approx 1$, indicate high linear dependence. Such redundancy is a major contributor to fragility in pairwise combinations. Employing correlated sensors adds little new unpredictability, which invalidates the simplistic additive model of entropy and provides diminishing returns on min-entropy. In other words, while individual modalities may appear moderately unpredictable, the correlations shown in Figure 2 can sharply constrain the overall joint entropy compared to what would be expected if they were independent variables.
4.3.1 Complexity Challenges
Computing the exact joint probability distribution and joint entropy of multiple sensors quickly becomes infeasible. Consider the following: let each of the $n$ sensor modalities be discretised into $b_{i}$ bins, then the joint distribution has $\prod_{i=1}^{n}b_{i}$ distinct states. This produces an enormous state space as $n$ and $b_{i}$ grow to the values considered in this paper. Applying the FD rule typically results in thousands of bins per modality, causing the number of joint bins to explode combinatorially. In addition, even before enumerating states, selecting which sensors to combine can itself involve $2^{n}-(n+1)$ subsets, skipping single-sensor subsets and the empty set. Preliminary experiments confirmed joint entropies could be computed directly for $n\leq 3$ modalities with a maximum of 1250 bins and fewer than 150K total samples from the Relay dataset. Reducing bin sizes can help, but this risks oversimplifying the distribution and artificially deflating entropy estimates, as we discussed §4.1. In our tests, we found that reducing bin counts reduced single-sensor entropy estimates by approximately 2–3 bits on average compared to those reported in Table 2. We therefore adopt an alternative strategy that trades some accuracy for tractability.
4.3.2 Chow-Liu Approximation
To handle these scaling issues, we adopt Chow–Liu trees [52] for approximating a high-dimensional joint distribution by a maximum-weight spanning tree, $\pi$, over the different sensor modalities. Each edge is weighted by the mutual information of the connected variables, ensuring the tree structure captures the dominant pairwise dependencies. This approach minimises the Kullback–Leibler divergence (Def. 2.2) between the true multivariate distribution and the resulting tree-based approximation as follows:
$p_{\pi}(x_{1},\dots,x_{n})=p(x_{r})\prod_{i\neq r}p(x_{i}\;|\;x_{\pi(x)})$ (9)
Where $\pi(i)$ denotes the parent of $X_{i}$ in the tree, and $r$ is the tree’s root node. Chow–Liu trees are acyclic, singly connected structures: each node has at most one parent where one can traverse the tree to accumulate probabilities between pairwise dependencies. The tree captures dominant pairwise dependencie and significantly reduces computation time compared to naïve enumeration of the full joint measurement space. The use of Chow-Liu trees was proposed by Buller and Kaufer [37] for estimating the entropy of multivariate data sources where the range of possible values is high, as is the case with sensor outputs considered in this work. In our Python implementation, we use the pgmpy [53] library’s TreeSearch module to this end; practically, for each sensor subset, we:
  1. 1.
    Discretise each sensor’s readings via Freedman–Diaconis binning.
  2. 2.
    Build a Chow–Liu tree from pairwise mutual information of sensors.
  3. 3.
    Traverse the tree to estimate max ($H_{0}$), Shannon ($H_{1}$), collision ($H_{2}$), and min-entropy ($H_{\infty}$) without enumerating the full state space.
Table 3: Average entropy results of Chow-Liu (CL) tree approximation against direct computation across all small sensor subsets ($n\in\{2,3\}$), giving the mean absolute error (MAE), averaged across all subsets, and the % mean relative error.
$H_{0}$ $H_{1}$ $_{2} HH_{\infty}$
Dataset $n$ Direct CL Direct CL Direct CL Direct CL MAE %
UCI-HAR 2 17.60 16.92 12.08 12.58 10.85 10.33 6.01 6.20 0.473 4.1
3 23.93 25.46 19.35 18.25 14.03 14.77 9.40 8.95 0.955 5.7
Relay 2 15.36 15.54 9.72 10.12 8.27 7.95 4.64 4.88 0.285 3.0
3 21.47 22.84 16.84 15.89 13.69 13.04 7.41 7.96 0.875 5.5
PerilZIS 2 20.33 22.18 14.77 14.00 11.40 12.18 7.21 8.52 1.013 7.6
3 30.82 32.55 21.28 20.08 16.66 17.65 11.46 12.06 1.258 6.2
SHL 2 24.63 26.25 18.76 18.20 14.17 14.28 9.50 8.92 0.693 4.1
3 36.71 39.05 28.44 26.83 22.46 21.14 12.44 13.25 1.488 4.9
We validated the approach against a ground-truth derived by computing the joint entropy through exhaustive enumeration. As this direct method is only computationally feasible for a small number of modalities, we performed the comparison on all two- and three-sensor subsets from each dataset. The results of this validation are shown in Table 3. Across all datasets, the estimates from the Chow-Liu tree closely track direct ground-truth calculations in these low-dimensional cases. The relative error is consistently low (3–8%), supporting its use of Chow-Liu trees for higher-dimensional combinations where direct computation is infeasible. [^5] We note the caveat that efficiently estimating the entropy of high-dimensionality sources remains an ongoing research challenge [37]. We must emphasise that our results are indicative estimations; quantifying errors beyond $n\geq 4$ is an open problem. Following this, we proceeded to estimating the entropy for all modalities in each dataset; our developed framework evaluates the joint entropy over all sensor combinations. The powerset of the sensor set is generated and processed in parallel using Python’s multiprocessing module. Processing all datasets took approximately 22 hours on our workstation with an Intel i7-6700K (8M cache, 4.20 GHz) and 32 GB RAM on Ubuntu 24.04.
4.3.3 Results
Tables 47 report the top 10 performing multi-sensor combinations ranked by min-entropy for each dataset. As expected, combining all sensors yields the highest $H_{0}$ (max-entropy) and often increases Shannon and collision entropy. However, min-entropy ($H_{\infty}$) remains stubbornly low. For instance, the complete set of sensors in SHL surpasses 80 bits of $H_{1}$ but saturates at only 21 bits of $H_{\infty}$. Interestingly, we find that omitting certain correlated sensors sometimes does not reduce min-entropy at all. For the Relay dataset in Table 5, the combination (Acc., Gyro., Light, Lin. Acc., Mag., Rot. Vec.) achieves $H_{\infty}=7.859$ bits, only slightly below the full set’s $8.092$ bits. Parallel findings arise in the PerilZIS and SHL datasets, where omitting a small number of sensors from the “All sensors” set has negligible impact on $H_{\infty}$. This pattern appears across datasets: additional modalities may raise $H_{0}$ and $H_{1}$ but barely move $H_{\infty}$.
(a) UCI-HAR
(b) PerilZIS
(c) Relay
(d) SHL
Figure 3: Average joint entropy vs. number of combined sensors modalities.
We plot the number of combined modalities for each dataset in Figure 3, showing the average (mean) joint entropy for different entropy measures. The data illustrates that, while $H$ values are tightly coupled initially, they diverge significantly as the number of joint modalities increases. All datasets demonstrated a trend of worst-case $H_{\infty}$ being significantly lower than other entropy values, often barely increasing relative to best-case metrics. For SHL and PerilZIS, for example, $H_{\infty}$ increases by approximately only 1–2 bits per modality, with $H_{0}$$H_{2}$ increasing far more significantly ($\Delta H_{0}\approx$ 10 bits and $\Delta H_{1}\approx$ 5 bits for every added modality). In Tables 47, we give the differences of $H_{1}$ and $H_{\infty}$ for the top combinations as $H_{1}-H_{\infty}$, quantifying the sharp decreases between Shannon and min-entropy as the number of combined modalities increases. We also use # to denote the number of modalities used in each row.
The results suggest that many sensors contribute redundant information when combined, exposing a fundamental limitation of multi-modal data in reality. Combining signals increases the apparent capacity for unpredictability, but redundancies between sensors means that the min-entropy from a large sensor ensemble is not substantially higher than a reduced subset thereof. This aligns with the emphasis of using min-entropy for worst-case guarantees in modern standards [25, 24]. To our knowledge, the extent of entropy collapse under correlated mobile modalities has not been elicited at this scale. Full combination results are available in our repository (see Footnote 3).
Table 4: Top 10 best-performing sensor combinations (UCI-HAR; in bits).
Modality # $H_{0}$ $H_{1}$ $_{2} HH_{\infty}$ $H_{1}-H_{\infty}$
All sensors 6 48.061 25.089 17.116 11.008 14.081
(Acc.{x,y,z}, Gyro.{y,z}) 5 39.403 22.835 15.999 10.113 12.722
(Acc.{x,y,z}, Gyro.{x,y}) 5 39.886 21.456 15.332 9.900 11.556
(Acc.{x,y,z}, Gyro.{x,z}) 5 40.222 21.470 15.040 9.411 12.059
(Acc.{y,z}, Gyro.{x,y,z}) 5 40.228 21.306 14.698 9.274 12.032
(Acc.{x,z}, Gyro.{x,y,z}) 5 40.293 21.260 14.575 9.154 12.106
(Acc.{x,y,z}, Gyro.y) 4 31.228 18.804 13.893 9.065 9.739
(Acc.{x,y}, Gyro.{x,y,z}) 5 40.273 21.155 14.208 8.721 12.434
(Acc.{x,y,z}, Gyro.z) 4 31.564 18.775 13.833 8.682 10.093
(Acc.{y,z}, Gyro.{y,z}) 4 31.570 18.583 13.400 8.386 10.197
Table 5: Top 10 best-performing sensor combinations (Relay; in bits).
Modality # $H_{0}$ $H_{1}$ $_{2} HH_{\infty}$ $H_{1}-H_{\infty}$
All sensors 7 45.794 25.036 15.725 8.092 16.944
(Acc., Gyro., Light, Lin. Acc., Mag., Rot. Vec.) 6 44.795 24.963 15.334 7.859 17.104
(Acc., Grav., Light, Lin. Acc., Mag., Rot. Vec.) 6 38.385 21.590 14.783 7.702 13.888
(Acc., Grav., Gyro., Light, Mag., Rot. Vec.) 6 37.026 21.001 14.553 7.702 13.299
(Grav., Gyro., Light, Lin. Acc., Mag., Rot. Vec.) 6 36.092 20.794 14.450 7.673 13.121
(Acc., Light, Lin. Acc., Mag., Rot. Vec.) 5 37.385 21.517 14.428 7.469 14.048
(Acc., Gyro., Light, Mag., Rot. Vec.) 5 36.026 20.924 14.270 7.465 13.459
(Gyro., Light, Lin. Acc., Mag., Rot. Vec.) 5 35.092 20.720 14.147 7.439 13.281
(Acc., Grav., Gyro., Light, Lin. Acc., Mag.) 6 41.094 22.794 14.437 7.338 15.456
(Acc., Grav., Light, Mag., Rot. Vec.) 5 29.617 17.555 13.244 7.312 10.243
Table 6: Top 10 best-performing sensor combinations (PerilZIS; in bits).
Modalities # $H_{0}$ $H_{1}$ $_{2} HH_{\infty}$ $H_{1}-H_{\infty}$
All sensors 13 83.662 36.998 33.682 23.926 13.072
(Acc.{x,y,z}, Light, Temp., Pres., Mag.{x,y,z}, Gyro.{x,y,z}) 11 76.907 34.258 31.510 23.835 10.423
(Acc.{y,z}, Light, Temp., Pres., Mag.{x,y,z}, Gyro.{x,y,z}) 10 72.737 34.246 31.509 23.835 10.411
(Acc.{x,z}, Light, Temp., Pres., Mag.{x,y,z}, Gyro.{x,y,z}) 10 72.820 34.243 31.508 23.829 10.414
(Acc.{x,y}, Light, Temp., Pres., Mag.{x,y,z}, Gyro.{x,y,z}) 10 72.737 34.229 31.501 22.945 11.284
(Acc.{x,y,z}, Light, Temp., Mag.{x,y,z}, Gyro.{x,y,z}) 10 71.384 32.586 30.190 22.945 9.641
(Acc.y, Light, Temp., Press., Mag.{x,y,z}, Gyro.{x,y,z}) 9 68.568 34.218 31.500 22.945 11.273
(Acc.x, Light, Temp., Pres., Mag.{x,y,z}, Gyro.{x,y,z}) 9 68.650 34.215 31.499 22.945 11.270
(Acc.{y,z}, Light, Temp., Mag.{x,y,z}, Gyro.{x,y,z}) 9 67.214 32.575 30.189 22.945 9.630
(Acc.{x,y,z}, Light, Temp., Pres., Mag.{x,z}, Gyro.{x,y,z}, Hum.) 11 76.618 33.816 31.035 21.915 11.901
Table 7: Top 10 best-performing sensor combinations (SHL; in bits).
Modality # $H_{0}$ $H_{1}$ $_{2} HH_{\infty}$ $H_{1}-H_{\infty}$
All sensors 22 158.601 82.301 39.320 21.289 61.012
(Acc.{x,y,z}, Gyro.{x,y,z}, Mag.{x,y,z}, Ori.{w,x,y,z}, Grav.{x,y,z}, LinAcc.{y,z}, Pres., Alt., Temp.) 20 148.995 78.624 39.276 21.289 57.335
(Acc.{x,y,z}, Gyro.{x,y,z}, Mag.{x,y,z}, Ori.{w,x,y,z}, Grav.{x,y,z}, LinAcc.{x,z}, Pres., Alt., Temp.) 20 148.759 78.477 39.272 21.289 57.188
(Acc.{x,y,z}, Gyro.{x,z}, Mag.{x,y,z}, Ori.{w,x,y,z}, Grav.{x,y,z}, LinAcc.{x,y,z}, Pres., Alt., Temp.) 21 148.587 78.380 39.249 21.289 57.091
(Acc.{x,y,z}, Gyro.{x,y,z}, Mag.{x,y,z}, Ori.{w,x,y,z}, Grav.{x,y,z}, LinAcc.{x,y}, Pres., Alt., Temp.) 20 148.952 78.135 39.235 21.289 56.846
(Acc.{x,y,z}, Gyro.{x,y,z}, Mag.{x,y,z}, Ori.{w,x,y,z}, Grav.{x,y,z}, LinAcc.{z}, Pres., Alt., Temp.) 19 139.153 74.800 39.175 21.289 53.511
(Acc.{x,y,z}, Gyro.{x,z}, Mag.{x,y,z}, Ori.{w,x,y,z}, Grav.{x,y,z}, LinAcc.{y,z}, Pres., Alt., Temp.) 20 138.981 74.703 39.128 21.289 53.414
(Acc.{x,y,z}, Gyro.{x,z}, Mag.{x,y,z}, Ori.{w,x,y,z}, Grav.{x,y,z}, LinAcc.{x,z}, Pres., Alt., Temp.) 20 138.745 74.556 39.117 21.289 53.267
(Acc.{x,y,z}, Gyro.{x,y,z}, Mag.{x,y,z}, Ori.{w,x,y,z}, Grav.{x,y,z}, LinAcc.{y}, Pres., Alt., Temp.) 19 139.346 74.458 39.100 21.289 53.169
(Acc.{x,y,z}, Gyro.{x,y,z}, Mag.{x,y,z}, Ori.{w,x,y,z}, Grav.{x,y,z}, LinAcc.{x}, Pres., Alt., Temp.) 19 139.109 74.311 39.087 21.289 53.022
4.3.4 Binning Sensitivity Analysis
As we discussed in §4.1 and §4.3.1, bin width is a key consideration in estimating entropy accurately and efficiently. Too few bins can lead to oversmoothing, where distinct features of the data distribution are lost, artificially deflating the entropy estimate. Conversely, too many bins can lead to sparsity and overfitting, where the analysis models the specific sample’s noise rather than the source’s true underlying distribution, inflating entropy. We performed a sensitivity analysis on the best-performing sensor combination (“All sensors”) for each of our four datasets. We re-calculated the joint entropy while varying the number of bins for each modality, starting from 5 bins and increasing up to 2048. Beyond that, the computational cost of processing the large state space became prohibitive. The results of this analysis are presented in Figure 4.
The graphs show a consistent trend across all datasets. For a low number of bins, all entropy measures are severely underestimated as expected. As the number of bins increases, the entropy rises sharply before tapering off and converging towards a stable plateau that approximates a logarithmic shape. (An exception is the min-entropy value for SHL, which converges after approximately 100 bins). Crucially, FD-selected bin counts fall within the stable region, supporting FD as a balanced choice between fidelity and tractability.
(a) UCI-HAR
(b) PerilZIS
(c) Relay
(d) SHL
Figure 4: Bin sensitivity analysis showing how entropy values vary by average bin width for the best-performing sensor combinations for each data set. Freedman–Diaconis (FD) values are also shown.
5 Evaluation
This section interprets our entropy findings for system design: we discuss the key takeaways, relate min-entropy to concrete guessing cost, discuss mitigations, and give recommendations for future research in the area.
5.1 Discussion
Prior studies have claimed that mobile sensors are suitable data sources for supporting security-critical applications. Much work has relied on model evaluation metrics as a proxy for evaluating such claims [7, 6, 21, 9], with a smaller subset using more established entropy metrics [36, 15]. Separate entropy analyses of sensors have demonstrated that individual sensors actually confer very little entropy, particularly under worst-case assumptions [28, 27, 26, 17]. Our results both challenge and extend these claims.
Single sensors yield low entropy. Building on prior work, we find that min-entropy often remains significantly below the Shannon estimates (e.g., some single-sensor readings yield only 1–3 bits). Our min-entropy estimates often remain significantly lower than their Shannon counterparts (e.g. some single-sensor readings yield only 1–3 bits of min-entropy). This exposes a major gap between average- and worst-case unpredictability across 25 widely used sensors, which is a critical distinction that existing schemes and evaluations do not capture at this scale.
Combining multiple sensors offers modest gains to worst-case min-entropy. While using multimodal sensors typically raises best- and average-case entropies ($H_{0}$, $H_{1}$), our experiments reveal that gains in worst-case min-entropy ($H_{\infty}$) are far smaller than one might expect. In many cases, $H_{\infty}$ plateaus, reflecting the dominance of a handful of high-probability joint outcomes. Even the most complex multimodal combinations in our datasets yield only modest worst-case entropy, well below the levels required by modern security standards. This also aligns with prior work showing that Shannon entropy can significantly overstate security [54, 55, 24, 25]. Hence, while mobile sensors can augment other authentication or key-generation processes, they do not suffice by themselves. Using multiple sensors may still improve usability and model accuracy, e.g. for coarse co-location checks, but they should not be treated as secure, unpredictable sources.
5.2 Translating Entropy Bounds Into Guessing Attacks
We sketch how low-entropy sensor data translates into concrete attack cost. Let $Z$ be the discretised joint sensor vector used by a scheme, and let $H_{\infty}(Z)$ denote its min-entropy. Ordering guesses by decreasing probability, the success probability after $q$ guesses satisfies
$\Pr[\text{success within }q]\;\leq\;q\cdot 2^{-H_{\infty}(Z)}.$ (10)
The expected number of guesses is $\approx 2^{H_{\infty}(Z)-1}$ [55].
Most systems accept similar, not exact, matches through learned decision boundaries or fuzzy matching [7, 6, 9, 21, 35, 20, 19]. Let us consider an abstract authentication function, $\mathrm{Auth}(z)\to\{0,1\}$, which accepts or rejects an input vector $z\in Z$. An adversary’s task becomes easier through a similarity-based scheme: they only need to generate a candidate $z^{\prime}$ that is ‘close enough’ to a high-probability vector $z$. This means the effective success probability for each guess is, in fact, higher than the bound in Eq. 10 suggests. The best-performing combination in our study yielded $H_{\infty}\approx 24$ bits, corresponding to a search space of $2^{24}$ ($\approx$16.7 million) possibilities. An adversary may pre-compute the most likely vectors and submit them to $\mathrm{Auth}$, requiring only $2^{23}$ guesses on average. This is well within the capabilities of a moderately resourced adversary—taking seconds to milliseconds on a modern PC to enumerate—and is orders of magnitude weaker than the $2^{128}$ expected of standard cryptographic keys [23]. Indeed, prior work has already demonstrated the feasibility of guessing biometric templates using this approach [42]. In the worst cases, such as the single-sensor readings where $H_{\infty}<8$ bits, the search space is trivial for an attacker to exhaust, requiring approximately $2^{7}=128$ guesses on average. While rate-limiting the authentication function can mitigate this risk, it does not solve the root problem. If each guess requires a live interaction, and the system enforces a limit of $r$ attempts per second, the expected time to success is scaled to $\approx 2^{H_{\infty}-1}/r$. Rate limits linearly stretch attack time but do not change the fundamental feasibility of guessing when $H_{\infty}(Z)$ is in the low-to-mid tens of bits.
5.3 What can be done?
Our results highlight a core tension in sensor-based security design. Proposals for pairing, co-location inference, or shared-secret agreement typically depend on the similarity of sensor signals across devices. Yet the very regularities that enable similarity seemingly stem from low-entropy structure. In effect, these systems trade cryptographic unpredictability for application-level utility. The open question is whether a scheme can ever achieve both meaningful entropy and the cross-device similarity it depends on. We close by offering several recommendations for researchers in this area.
Evaluate the threat of adversarial samples
A persistent feature in the literature is the lack of rigorous analysis of min-entropy (and induced guesswork) under realistic acceptance regions and attacker priors [49, 7, 9, 10, 21, 8, 5, 6]. We urge the community to consider the resilience of proposals to adversarial samples, particularly when using ‘fuzzy’ matching methods such as supervised learning models, and to translate those estimates to concrete costs, such as expected guesses.
Quantify effects of deliberate interaction and induced phenomena
User actions such as prescribed shakes or gestures, as well as externally induced signals, can inject variability and modestly raise entropy by several bits [26, 28]. For example, altering the device’s environment to produce distinctive electromagnetic signatures may be detectable by onboard sensors. A key gap is to quantify how much $\Delta H_{\infty}$ such interventions can actually deliver across modalities. Our expectation is that any gains remain modest, but this trade-off between added entropy and practical feasibility remains an under-explored area of research.
Decouple sensor similarity from unpredictability
Where applications need similarity evidence, such as for device authentication, one should treat that evidence as weak, predictable inputs. The derivation of shared secrets must come from a component designed for unpredictability; any sensor contribution should be considered, at most, as a small additive input to an entropy pool (see [38, 40]). Cryptographic extractors (e.g. Von Neumann and later constructions [56, 57, 58]) or decorrelation techniques (e.g. PCA) may be considered to convert low-quality sensor noise into higher-entropy sources. However, while extractors reduce bias and correlation, they cannot increase intrinsic min-entropy. That is, if sensors offer $k$ bits of $H_{\infty}$, post-processing can yield (at best) an unbiased string of length $\approx k$, and no more. Extractors improve the quality of usable randomness when sufficient $H_{\infty}$ already exists; they do not increase the quantity in a way that raises brute-force cost. Future work in the area should focus on how entropy can be meaningfully increased while retaining application utility.
5.4 Scope and Limitations
While we analyse four large datasets, they do not fully capture contexts all possible contexts. Some work has suggested that performing dedicated movements, such as gestures, can increase the amount of usable entropy from motion sensors in the region of approximately 5–6 bits [28, 26]. Our datasets do not cover such dedicated movements; it is possible that the reported results are an underestimation of entropy for sensors which are deliberately perturbed as an entropy-generating action. Conversely, we do not examine entropy-reducing threats, such as sensor spoofing [10], hardware attacks [44], or cross-device correlation attacks [45]. It is concievable that these threats may reduce the entropy of sensors even further. [^6] By way of example, we could consider a voltage-based fault injection attack that causes nondeterministic errors in the output of individual sensors, affecting the entropy of their output values.
We also estimate unconditional entropies such that all measurements are independent and identically distributed (i.i.d.). This is an optimistic assumption, as sensor data is inherently time-correlated; a sensor’s next reading is often highly dependent on its previous readings. For Shannon entropy and for the conditional Rényi notions standard in cryptography (e.g. min-entropy), the conditional entropy is always less than or equal to the unconditional entropy; that is, $H_{\alpha}(X\,|\,Y)\leq H_{\alpha}(X)$ (see [59] for additional properties). Thus, the unconditional entropy values we calculate in this work act as an upper bound. We may wish to condition $X$ on a sequence of previously observed measurements; however, efficiently quantifying $H_{\alpha}(X_{t}\,|\,X_{<t})$ for high-dimensional sensor streams remains non-trivial. Nevertheless, we posit that this would only serve to decrease entropy, strengthening our core conclusion that mobile sensors are insecure data sources for security applications.
6 Conclusion
This paper presented the largest study to date on the entropy of mobile sensors across four public datasets (UCI-HAR, SHL, Relay, and PerilZIS) and 25 modalities. Our results expose a gap between the perceived and actual strength of sensor data for security: even when max- or Shannon entropies appear high, worst-case min-entropy often collapses to insecure levels. Modalities that appear strong under average-case metrics, which can mask inherent biases, are insecure under worst-case measures advocated in modern security standards. Consequently, this creates the opportunity for adversaries to guess the most probable values with minimal effort.
Our findings also challenge the common practice that model evaluation metrics (e.g. accuracy or EER) suffice to demonstrate that sensor-based security applications are, indeed, secure. The vulnerability of mobile sensors to biased distributions and inter-sensor correlations significantly undermines their effectiveness as reliable sources of unpredictability. The effectiveness of countermeasures remains an open and pressing research challenge without compromising security or utility. Our hope is that the methodologies and insights presented here will encourage the security community to adopt more rigorous evaluation strategies for sensor-based techniques, paving the way for safer and more robust designs.
In future work, we consider that a dynamic analysis is important to assess the stationarity issues with sensor data, where entropy varies according to user behavior or environment. A user study could yield empirical bounds on how finely humans can intentionally manipulate motion sensors, revealing more realistic limits to sensor-based randomness in real-world scenarios. Our binning decision was a statistical one, rather than one that reflects human usage; it is possible that real-world usage may reduce the resolution of useful sensor data, reducing entropy even further. Overall, while sensor-based data can reach limited levels of entropy under favourable conditions, the road to making such sources systematically secure and robust is long. The key takeaway is that substantial work is needed before mobile sensors can be considered appropriate for security-critical applications.
Acknowledgments
Carlton Shepherd has received funding from the UK EPSRC ‘Chameleon’ project (EP/Y030168/1). The authors would like to thank Darren Hurley-Smith for insightful conversations on the topic.